A Complete Bibliography of ACM Transactions on the Web (TWEB)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

23 August 2024 Version 1.50

Title word cross-reference

1 [PYWY23].

2 [PYWY24]. 2.0 [KL10].

4chan [MZBD20].

7 [UÁM17].

AA [BZR^{+}21]. Abstract [AZ19]. Academic [WLT^{+}21]. Accelerate [CJQ^{+}21]. Acceleration [HKH^{+}16]. Access [KL17a, PMOB11, SB24, SRG^{+}22, CSCB07, DFJ^{+}12]. Accessibility [HR13]. ACConv [PMOB11]. Account [KMV15, MNB^{+}24]. Accuracy [HHBT20].

Across [SQL^{+}23]. Actions [LZKN22, RDM09]. Active [CZZ15]. Activities [WPC^{+}22]. Activity [CLR^{+}19, TCM^{+}18, WHGS16, ZJZ^{+}21]. Actually [DZS^{+}16]. Adaptation [SDG^{+}23]. Adaptive [GKS^{+}08, JWJ^{+}18, MTDF18, WWC^{+}24, YYL^{+}23, ZLG^{+}24]. addressing [AMND^{+}08]. Administrators [DLMI16]. adopting [VGA13]. Adoption [LLA^{+}24, RHLC17]. Ads [AY10]. Ads-portal [AY10]. Advanced [PYWY23, PYWY24, SDN08]. Adversarial [ZZL^{+}24, ND08]. Advertising [PAAC17, LHJL13, VGA13]. Aesthetics [UÁM17]. African [FTF^{+}18]. against [JTLM24, ZZL^{+}24]. aggregate [UT11]. Aggregations [ZYZ16]. Aggression [PVK22]. agreement [BKJ13]. Ajax [MvDL12]. Ajax-Based [MvDL12]. AjaxScope [KL10]. Alert [SGJC20]. algebra [YB08]. Algorithms [AH23, BHMW11, Hog17, PSP23, PYWY23, PYWY24, CCFF11, SZG11, YZL07]. Aligned [TBBI18]. Alignment [SDG^{+}23]. alike [DAA13]. Among [ABLW19, DLMI16]. Analysing [FTF^{+}18]. Analysis [ACC24, AH23, AALG22, BSW23, CRB18, EYH16, GFTC19, MHG^{+}20, MMMD16, MvDL12, OSB22, PVS19, RSC^{+}21, SSK^{+}17, SGJC20, SAS24, TWH14, WPC^{+}22, AB08, BCD^{+}08, LV13, LSC^{+}08, SMB^{+}07]. Analysis-driven [OSB22]. Analytic [UPS^{+}07]. analytics [AMND^{+}08]. Analyzing [BCGL17, CASN13, MZBD20, RHLC17, SCS^{+}14, WSL^{+}19, WCZ15]. Android [MHG^{+}20]. annotation [BEP^{+}08, NKTP13]. Annotations [WSL^{+}19]. Anomalies [DWXL23]. Anomalous [WQG^{+}21]. Anonymization [SLM13]. Answering [SB16]. Anxiety [MP23]. Anycast [ALR^{+}11]. AOI [EYH20]. API [JTLM24, WCZ15]. APIs [BZA24, BDM17]. App [DV18, LLA^{+}24, SLP^{+}19, MHG^{+}20]. Application [BBBF14, DBZ^{+}12, DGS17, JJM20, WLP^{+}23, SIYL08]. application-level [SIYL08]. Application-Specific [DBZ^{+}12]. Applications [ACC24, KL10, MvDL12, PYWY23, PYWY24, RHLC17, VP11, WDD15, WdSdO23, CM12, DFJ^{+}12, DWS^{+}12, FCBC10, UPS^{+}07, ZCL^{+}10, CGM14, DJBO14]. Approach [BZR^{+}21, BCC14, CTC^{+}15, CMP15, DJBO14, JWJ^{+}18, JXCX21, OACU13, RSC^{+}21, SZSA15, SFJ^{+}23, WZZ^{+}16, XZ17, ARN12, FCBC10, LJP^{+}13, MSBB10, QBC13, SIYL08]. Approaches [EYH20, PAC^{+}21, SB16, ZWZL15, ETT08]. Apps [MHG^{+}20, SSK^{+}17]. Arbiters [DLMI16]. Architecture [ALR^{+}11, YYL^{+}23, LMJ10, SRRG07]. architectures [RCS^{+}08, ZHD07]. Archive [JKWN24]. Archived [BKNW23]. Ascertain [UBL^{+}24]. Aspect [SZSA15, LCZ^{+}20]. Aspect-Based [SZSA15]. Assessing [BKJ13]. Assistant [CRPLM11]. Association [WLCG20]. Association-rule-based [WLCG20]. Assortativity [SÖ23]. Attacks [JTLM24, MNB^{+}24, MKH22, JBB^{+}09, SIYL08]. Attendance [ZJZ^{+}21]. Attention [ALS23, CLZ^{+}22, PWH16, VVB21]. Attribute [FLT15]. Attributed [LMLC24, WLT^{+}21]. auctions [RDJS07]. augmentation [DAA13]. Augmented [DA15]. Authentication [AALG22, CTC^{+}15]. Auto [ZCC^{+}23]. Auto-Encoders [ZCC^{+}23]. Autocompletion [WMS^{+}16]. Automated [DWS^{+}12, SCB17, WSL^{+}19]. Automatic [BEP^{+}08, CR20, CB20, EYH20, MMH13, SZG11]. Automatically [JKWN24]. average [WOHM08]. AVX2 [ML18]. Aware [AYJS24, ABO^{+}16, BCF16, OAU11, CB20, ETT08, LJP^{+}13, LTS^{+}24, MWHW24, QBC13, STZL20, WZZ^{+}16, ZLW^{+}22].

Backup [RH19]. Balance [QA14]. Bandit [WOM22]. BanditProp [WOM22]. Base64 [ML18]. Based [AYJS24, AADP19, BHMW11, BHW13, BCC14, BBBF14, CRB18, CAOU12, DJBO14, FLT15, GAC^{+}11, GZYS16, HGPS11, JWJ^{+}18, KKSS17, LCZ^{+}20, MA14, MvDL12, MAY^{+}11, PCCS23, PKT17, RHLC17, SZSA15, TBBI18, UÁM17, WPB13, WGF24, ZWML14, ZWH^{+}24, ZCC^{+}23, ABLW19, AADS13, BKJ13, BEP^{+}08, CKJA13, CJQ^{+}21, CP09, LHJL13, LLM13, MWHW24, MFB21, PAL18, SFJ^{+}23, VVCD13, WQG^{+}21, WLCG20, YL08, YZD^{+}24, ZJZ^{+}21, ZCL^{+}10, ZZM^{+}11, ZLL^{+}23, ACC08]. Bases [ZTC11]. Bayesian [GZYS16, SHHS17, WML23].

Behavior [CBB18, DZS^{+}16, ESTD24, KL10, LTS^{+}24, RH19, TWH14, WZT^{+}17, CSCB07, LV13, ZZW^{+}24]. Behavior-aware [LTS^{+}24]. Behavioral [SAS24]. BehaviorNet [LTS^{+}24]. Behaviors [XLC20]. Behavioural [GFTC19]. Behind [JBWR20]. Being [CCZP24]. Benchmarking [AST19]. BERT [OSB22]. Best [ABO^{+}16, EYH20]. Better [YZZ^{+}24]. Beyond [CGM14, WSPZ12, LLWL09]. Bias [RDTT24, WH14]. biases [JGTF10, WH13]. BibSonomy [DZS^{+}16]. bid [RDJS07]. bidding [RDJS07]. billion [LLWL09]. Bitcoin [WPC^{+}22]. Bitemporal [FGH^{+}16]. Blank [Hog17]. Blitz [ESTD24]. block [CJQ^{+}21]. Blog [JKH^{+}12, WYY^{+}15, APV10]. blogosphere [MMB^{+}12]. BNoteHelper [YZD^{+}24]. Bot [CLR^{+}19, YYL^{+}23]. Both [EYH20, ZWML14]. Bots [GFTC19]. bottom [QBC13]. bottom-up [QBC13]. bounds [VKY10]. BPEL [SZSA15]. Brand [CR20, RSC^{+}21, SW11]. Breast [PWH16]. Bridging [CCZP24]. Bring [KSG^{+}22]. Bringing [PAAC17, XLC20]. Browse [CCC22, PGAW23]. Browser [AALG22, DA15, LBBA20, WAP19, WdSdO23]. Browser-Side [DA15]. Browsers [GTK11, JBB^{+}09]. BrowserShield [RDW^{+}07]. Browsing [TWH14, VDM^{+}18, XLH^{+}09, DWC12, LV13]. BUbiNG [BMSV18]. Building [TKS11]. business [LMJ10, PSBY10]. Butterflies [SÖ23]. Buyer [Fra16]. Buyer-Friendly [Fra16].

Cache [CAOU12, CJQ^{+}21, SRRG07]. Cache-Based [CAOU12]. Cache-block-based [CJQ^{+}21]. Caching [CJQ^{+}21, DV18, OAU11, OACU13, BYGJ^{+}08]. Camera [SW11]. Campaigns [ZLZL16]. Cancer [PWH16]. Canonical [Hog17]. Captions [WH13]. Caring [DAA13]. Cascading [RHLC17]. Case [LLA^{+}24]. Cashtag [CLR^{+}19]. Castle [MHG^{+}20]. Categorizing [PAC^{+}21]. Category [MWHW24]. Category-aware [MWHW24]. Causality [WGF24]. CBPCS [CJQ^{+}21]. CDN [ALR^{+}11]. Censorship [GCND^{+}15]. Center [SGJC20, DWS^{+}12]. centralities [YL08]. Centrality [LYZ23]. Centric [BSW23, CMP15, TTHS19]. Certification [AADP19, AADS13]. Challenges [DA15, GAC^{+}11]. Chance [OACU13]. Change [WQG^{+}21]. Change-based [WQG^{+}21]. Changes [MvDL12]. Changing [LLSL18]. Channel [LLW^{+}23, MKH22, QWW^{+}24]. Characteristics [SSK^{+}17, WST11, ZLD^{+}21]. Characterization [JKH^{+}12, PGAW23, VDM^{+}18]. Characterizing [CBB18, GAC^{+}11, GCND^{+}15, MAY^{+}11]. ChatGPT [LFAH24]. Chief [Whi21]. choreographies [MPvdA^{+}10]. Cities [SQL^{+}23]. Claims [SKGY14]. Classification [ALS23, BHMW11, BHW13, GCMG15, HY23, HHBT20, RDW^{+}16, SHLS24, SB16, WLP^{+}23, WAP19, WdSdO23, DGP09]. Classifying [GBF^{+}09, PMS^{+}24]. CLHHN [MWHW24]. ClickRank [ZM12]. Clickstream [WZT^{+}17]. Client [CTC^{+}15, KL10, SIYL08]. Client-Side [KL10]. client-transparent [SIYL08]. Closeness [LYZ23]. Clustering [EYH16, KH15, KV11, WWC^{+}24]. Clustering-Driven [KV11]. Coding [WML23]. Coefficients [KH15]. COIP [BCF16]. COIP-Continuous [BCF16]. ColBERT [WMTO23]. ColBERT-PRF [WMTO23]. Cold [GCH^{+}21, MPB20]. Cold-start [GCH^{+}21, MPB20]. Collaboration [Lee15, WHGS16, WLT^{+}21]. Collaborative [BCC14, LLW12, WWW^{+}17, CCFF11, MKR07, RHS09, VGA13]. Collections [WBdR12, CMRV10].

Collective [DLMI16]. Collusive [XZ17]. Combating [KEG^{+}08, ZWML14]. Combined [SHLS24]. Combining [MDG19, WdSdO23]. Comment [HGC^{+}18, SCS^{+}14]. Comments [LFAH24, SCS^{+}14, WYY^{+}15]. Commerce [PKT17, ZWZL15, Jan07]. Communal [UC22]. Communication [RUK19]. Communities [Hua13, Lee15, DGP09, YL08]. Community [GLY^{+}24, KMSB24, LMLC24, SB16]. Community-enhanced [KMSB24]. Compact [ÁL16, CN10]. comparative [Jan07, SMB^{+}07]. Comparing [CPX14, SHHS17, ZBG^{+}15]. Comparison [AST19, CCFF11]. compensations [SDN08]. Completeness [DNPR18]. Completion [LZX^{+}24]. complex [DWS^{+}12]. Complexities [WHS13]. Composite [AADP19]. Composition [BAP13, CBB17, CB20, MFB21, WPB13, ZLKL19, ZYZ16, ARN12]. Compositions [ELM16, KVSH23, SZSA15]. Comprehensive [BHMW11, BHW13, SB16]. compressed [VKY10]. Compression [MKH22, NRS^{+}22]. Computation [ZWZL15]. Computational [AKZ20]. Computer [CCC22]. Computing [CLB19a, ZTC11, DK08, WLCG20]. Concept [GWL^{+}23]. Concepts [AL24]. Conceptual [SDC14]. confidentiality [PSBY10]. Confirmation [RDTT24]. conflict [AMND^{+}08]. Conformance [SFJ^{+}23]. Congruence [SW11]. Connections [WST11]. connectors [MKR07]. Consensus [SXM^{+}16, RHS09]. Consent [JTVM22]. Considering [GLY^{+}24]. Consistency [RSC^{+}21]. Constrained [PMGO18, SMRM07]. Constraint [SFJ^{+}23]. Constraint-based [SFJ^{+}23]. constraints [ARN12, YZL07]. Constructing [CPX14, SXY^{+}23]. Construction [ELM16]. consumers [DAA13]. Consumption [DV18]. Content [BBBF14, CRB18, RDW^{+}16, RSC^{+}21, WH14, ZBG^{+}15, CM12, LLM13]. Content-Based [BBBF14, LLM13]. Context [AYJS24, Fra16, ZLW^{+}22, ZM12, LJP^{+}13, LLM13]. Context-Aware [AYJS24, ZLW^{+}22, LJP^{+}13]. Contextual [ZWZL15, LHJL13, VGA13]. Continuous [BCF16, CZ21]. contracting [CP09]. Contrastive [WZJ^{+}24, WML23]. Contribution [PSP23]. Control [BAP13, PMOB11, PAAC17, SRG^{+}22, DFJ^{+}12]. Control-Flow [BAP13]. Controlled [DGS17]. Conversational [PMOB11, WCF^{+}23]. Conversations [AYJS24]. Conversion [DBZ^{+}12]. Convolutional [CLZ^{+}22, SDSS24, ZZW^{+}24]. Cookies [TM09]. Coordinating [KWLA13]. Corpora [JKWN24]. Corporate [FG18]. Correctness [ETT08]. Correctness-aware [ETT08]. Correlation [WGF24]. Cost [OAU11]. Cost-Aware [OAU11]. Counterfeit [CR20]. crawl [BYKS09]. Crawling [BMSV18, DJBO14, MvDL12, VCK14]. Crawls [CSLL18, SMB^{+}07]. Creators [SAS24]. Credibility [SDSS24]. Credit [GZC^{+}16]. critical [CCH^{+}22]. Cross [GCH^{+}18, GCH^{+}21, WAP19, WdSdO23, ZZL^{+}24]. Cross-Browser [WAP19, WdSdO23]. Cross-Platform [WAP19]. Cross-Site [GCH^{+}18, GCH^{+}21, ZZL^{+}24]. Crossing [ZWH^{+}24]. Crowds [LCK^{+}12]. Crowdsourced [UBL^{+}24]. Crowdsourcing [ABO^{+}16, MPB20, TDKC15, WSL^{+}19]. CSS [MMMD16]. CSS-Sprite [MMMD16]. current [CCFF11]. Custom [SDC14, TDKC15]. Customers [MNB^{+}24]. Cyber [ABLW19]. Cyberaggression [CLB^{+}19b]. Cyberbullying [CLB^{+}19b, CZ21]. Cybersecurity [SGJC20].

DanMu [HGC^{+}18]. Data [ABO^{+}16, AKJ^{+}18, AZ19, BGNV10, DNPR18, EYH20, GCND^{+}15, GCMG15, GZC^{+}16, KI23, PMGO18, RP17, TTHS19, Tho14, ZJZ^{+}21, DFJ^{+}12, DWS^{+}12, QBC13, ZCL^{+}10, FPG15]. data-intensive [DFJ^{+}12]. Data-Types [AZ19]. Databases [PPPS18]. Datasets [LCLQ19, RKOD22, TBBI18, WCZ15]. Day [LZKN22]. Debated [RDTT24]. Decentralized [BAP13, KVSH23]. decisions [ACC08]. Declarative [MPvdA^{+}10]. Decoding [ML18, SMB^{+}07, XVWK23]. Decompositions [SSPÇ17]. Decoupled [CYL^{+}21]. Deep [CCL^{+}23, SDG^{+}23, TC20, WWC^{+}24, WWN^{+}20, ZC23, BKJ13]. Defending [ZZL^{+}24]. Defense [FKW^{+}21]. DeFi [KVSH23]. Defined [PPPS18]. definitions [BEP^{+}08]. DeLink [ZZL^{+}24]. Deltas [ZTC11]. demand [SRRG07]. Demonstration [TKS11]. denial [SIYL08]. Dense [SSPÇ17, WMTO23, DGP09]. Dependent [WST11]. Deployment [CRB18, TM09]. Depression [MP23]. depth [SAS24]. DescribeX [CMRV10]. Descriptors [UÁM17]. Design [BYGJ^{+}08, WdSdO23, ZLKL19]. Designing [GTK11]. destinations [WBC08]. Detecting [ABLW19, CLB^{+}19b, CASN13, LFAH24, LSC^{+}08, ZLZL16]. Detection [AKZ20, ALS23, CR20, CZ21, DWXL23, EYH20, GLY^{+}24, Lee15, LLW^{+}23, OSB22, SSK^{+}17, SXY^{+}23, WAP19, XZ17, YYL^{+}23, ZJ20, AMND^{+}08, BCD^{+}08, BLW13]. Deterministic [BGNV10]. Developer [CCH^{+}22]. Development [CMP15, KPED14, SDC14, VP11]. Device [DWXL23, VDM^{+}18]. Device-level [DWXL23]. diagnostic [WH13]. Different [SQL^{+}23, BYKS09]. Differentiation [ZWML14]. Difficulty [Tho14]. Diffusion [Gae18, JKS23, MSP^{+}17, PVK22]. Digital [KL17b]. Dimension [KI23]. Dimensional [BDM17]. Directed [CYL^{+}21, SMRM07]. directory [LHJL13]. Dis [RH19]. Disasters [RGGG18]. Discourse [MZBD20]. Discovering [ABO^{+}16, WMS^{+}16, YL08]. Discovery [APV10, MDG19, PT09, WJH13]. Discriminating [LFAH24]. Discussions [BYJ^{+}21]. Disentangling [KVSH23]. Disk [MA14]. Disk-Based [MA14]. Disorder [ALS23]. displays [XLH^{+}09]. Dissemination [PSP23]. Distance [KKSS17, ZLW^{+}22, VKY10]. Distillation [WST11]. Distributed [MK12, SLL^{+}15, WLCG20, CKJA13, LMJ10]. Distributed-computing [WLCG20]. Distributions [RDW^{+}16, WWC^{+}24]. Distrust [QA14, ZWML14, VVCD13]. Diversification [NAS16]. Diversionary [WYY^{+}15]. Diversity [ZLG^{+}24]. DNS [JBB^{+}09]. Do [BYKS09, CCC22, DZS^{+}16]. Document [WMTO23]. Does [PMGO18]. DOM [BLW13, WAP19]. Domain [CSLL18, SDC14, SDG^{+}23, WPB13]. Domain-Specific [SDC14]. domains [AY10]. Double [CLZ^{+}22]. Download [JBWR20, Uzu24]. Drive [JBWR20]. Drive-by [JBWR20]. Driven [EYH20, KV11, VP11, BLW13, CM12, FCBC10, OSB22, RDW^{+}07, SFJ^{+}23, ZHD07]. Dual [QWW^{+}24]. Dual-channel [QWW^{+}24]. Duplicate [SHLS24]. During [CCZP24, UC22, RDTT24, RGGG18]. DUST [BYKS09]. Dynamic [AL24, CZ21, JJM20, KMSB24, LTS^{+}24, MvDL12, OACU13, WQG^{+}21, WML23, XJZL24, ZLW^{+}22, RDW^{+}07, RCS^{+}08]. Dynamics [LLSL18, MSP^{+}17, WHGS16, LAH07, LSC^{+}08].

E-Commerce [PKT17, Jan07, ZWZL15]. Early [DQSZ19, ZJ20]. eBay [RDJS07]. Ecommerce [XLWS17]. Ecommerce-Reputation-Escalation-as-a-Service [XLWS17]. Ecosystem [FTF^{+}18, PVS19, XVWK23]. Edge [PPPS18]. Edge-Labeled [PPPS18]. Editor [Whi21]. Editor-in-Chief [Whi21]. Editorial [Ano15, AIN12, Dav18b]. Effective [WOM22, WGF24]. effectiveness [Jan07]. effects [RDJS07]. Efficient [BYG11, NAS16, PBSO14, SLM13, YZL07, YMZ19, ZWZL15, ARN12, CKJA13]. Effort [KL17a, Lee15]. Elastic [FLT15]. Electrical [LLW^{+}23]. Element [VVB21]. Elements [WAP19]. Elimination [CCH^{+}22]. Embedded [KPED14]. Embedding [CYL^{+}21, JWJ^{+}18, WLT^{+}21]. Embeddings [ABLW19, MA22, OSB22]. Emergence [RHS09]. Emerging [HGC^{+}18, CSCB07]. Emotions [JBWR20]. Empirical [ACC24, AALG22, FG18, MHG^{+}20, MNB^{+}24, RHLC17, WPC^{+}22, XLWS17, MSBB10, RDM09, WOHM08]. Enacting [TDKC15]. Encoders [ZCC^{+}23]. Encoding [ML18]. End [CMP15, ZLKL19, ARN12, DAA13, YZL07]. end-to-end [ARN12, YZL07]. End-User [CMP15, ZLKL19, DAA13]. Enemy [ABLW19]. Energy [DV18]. Engagement [DGS17]. Engine [BWLK10, BYG11, DGS17, MP23, SHKK14, BYGJ^{+}08]. Engineering [FCBC10]. Engines [CAOU12, DDB^{+}14, OAU11, OACU13]. enhance [WBC08]. Enhanced [ZZW^{+}24, CCL^{+}23, KMSB24]. Enhancing [VVCD13, WCF^{+}23, Coo08]. Enrich [ZM12]. enterprise [GKS^{+}08]. Entity [MDG19, TTHS19, XLC20, WJH13]. Entity-Centric [TTHS19]. Enumeration [MNB^{+}24]. environment [SDN08]. Environmental [AL24]. Environments [ZWZL15]. Envisioning [SB24]. Epidemic [RKOD22]. Episodic [PWH16]. Equivalent [Hog17]. Era [KL17b]. Escalation [XLWS17]. Estimating [Hua13, KH15, SXM^{+}16]. Estimation [BCF16, SDSS24]. ethics [AB08]. Evaluating [FG18, PKT17, WQG^{+}21].

Evaluation [CLU16, KEG^{+}08]. Evaluations [RDW^{+}16]. Event [FGH^{+}16, LLW^{+}23, SLM13, ZJZ^{+}21]. Event-based [ZJZ^{+}21]. Events [FGH^{+}16, MJ17]. Evidence [GZC^{+}16, GZYS16]. Evidence-Based [GZYS16]. evolution [RCS^{+}08]. Evolutionary [BZA24]. evolving [NRS^{+}22]. Exclusive [MKH22]. Executable [ELM16]. Execution [CJQ^{+}21, JJM20, LMJ10]. Exemplars [JKWN24]. Exif2Vec [UBL^{+}24]. EXIP [KPED14]. Expansion [WBdR12, XLC20]. experience [BKNW23]. Experimental [CR20]. Experiments [DGS17]. Expert [ZBG^{+}15]. Experts [WPB13]. Explain [Tho14]. Explainable [WGF24]. Explaining [SÖ23]. explicit [VVCD13]. Exploiting [CASN13, LLSL18, SLP^{+}19, WBdR12]. Exploration [MP23, DWC12]. Exploratory [GZC^{+}16]. Exploring [BCGL17, CMRV10, FTF^{+}18, HGC^{+}18, ZJZ^{+}21]. Expressions [BGNV10]. Extended [QA14, SRG^{+}22]. External [ÁL16, WBdR12, UT11]. Extracting [MMB^{+}12, PMS^{+}24, RGGG18, RN09]. Extraction [DGP09, FGH^{+}16]. Extractor [AST19]. Extractors [AST19]. Eye [EYH16, EYH20].

Face [SXY^{+}23]. Facebook [CHC13].

Factorizing [LZKN22]. Failure [WdSdO23]. Fair [BPG^{+}22]. Fake [RKOD22, DQSZ19]. Fast [CBB17, CN10, LCLQ19, MA14, VKY10]. Faster [ML18]. Feature [VVB21, WdSdO23, NKTP13]. feature-word-topic [NKTP13]. Features [BHMW11, GLY^{+}24, MDG19]. Feedback [PKT17, SKGY14, WMTO23, ZXC^{+}24]. Feedback-Based [PKT17]. Few [LZX^{+}24]. Few-Shot [LZX^{+}24]. Figurative [AKZ20]. Filtering [BCC14, CCFF11, MK12, MKR07, RDW^{+}07, VGA13]. Filters [SLP^{+}19]. Finance [KVSH23]. Financial [CLU16]. Fine [LTS^{+}24]. Fine-grained [LTS^{+}24]. Fingerprinting [LBBA20]. Fingerprints [AALG22]. FinTech [CHT^{+}23]. Fisher [WWC^{+}24]. Five [AST19]. Flares [UC22]. Flexible [ELM16, FPWC22, SDN08]. Flickr [RN09, SW11]. Flow [BAP13, HGPS11]. Focus [WDD15]. Focused [VCK14]. Fona [WDD15]. Footprint [GZC^{+}16]. Forecasting [AL24]. Forgery [SXY^{+}23]. Form [WPB13]. Form-Based [WPB13]. Formal [FPG15, KMV15]. formerly [CCZP24]. Forms [Hog17]. Formulation [AH23]. forums [JGTF10]. Foundations [QA14]. FoXtrot [MK12]. Framework [AMH^{+}21, BKNW23, BDM17, KV11, KPED14, QWW^{+}24, SHLS24, UBL^{+}24, WLCG20, XLC20, YB08, ZZL^{+}24, ZLKL19, ACC08, CP09, WJH13]. Frameworks [JCD^{+}21]. Fraud [XZ17]. Fraudulent [SAS24]. Free [PVS19, VVB21]. Free-viewing [VVB21]. Friendly [Fra16]. friends [ZZM^{+}11]. Friendship [ABS^{+}12]. Ful [FGH^{+}16]. functional [ETT08]. FusE [TTHS19]. Fusion [TTHS19, WCF^{+}23, ZC23]. Fuzzing [BZA24]. Fuzzy [ZS24].

Game [ESTD24, SRG^{+}22]. Gated [ZC23]. Generalized [WQG^{+}21]. Generating [AKJ^{+}18]. Generation [JCD^{+}21, LCLQ19, YZD^{+}24]. Genetic [MZBD20]. Georeferenced [DDB^{+}14]. Geosocial [KKSS17]. Gifting [RHLC17]. global [YL08]. GNN [YYL^{+}23]. GNNs [LLW^{+}23]. Goal [ZLD^{+}21]. Good [SB24]. Governance [AL24]. GPS [ZCL^{+}10]. grained [LTS^{+}24]. Granularity [KI23]. Graph [AYJS24, ALS23, BCGL17, CN10, FPG15, JKS23, JWJ^{+}18, LZX^{+}24, PPPS18, PYWY23, PYWY24, SDSS24, SW11, WLP^{+}23, WZJ^{+}24, WWC^{+}24, WGF24, XJZL24, ZLG^{+}24, ZZW^{+}24, ZCC^{+}23, ZXC^{+}24, DGP09]. Graph-Based [AYJS24]. Graphlets [JXCX21]. GraphQL [BZA24]. Graphs [Hog17, LYZ23, QWW^{+}24, SXY^{+}23, WSPZ12, YZZ^{+}24, ZWH^{+}24]. Group [SDG^{+}23]. GroupAligner [SDG^{+}23]. Grouping [HHBT20].

Habits [VDM^{+}18]. Hate [ABLW19]. Hateful [LFAH24]. healing [AMH^{+}21]. Health [WH14]. Healthy [SB24]. Helpfulness [PCCS23]. Here [PAL18]. Heterogeneous [AL23, GWL^{+}23, MWHW24, VBM17, WLP^{+}23, ZLG^{+}24, ZZW^{+}24, ZWH^{+}24]. Hidden [BCGL17]. hierarchically [XLH^{+}09]. Hierarchy [SSPÇ17]. High [CLB19a, CCFF11, ETT08]. high-level [ETT08]. High-performance [CLB19a, CCFF11]. Historic [BWLK10]. Historical [LZKN22]. Histories [PWH16]. history [ZZM^{+}11]. Hoaxes [ZJ20]. Home [CCC22, DWXL23, KWLA13]. Homepage [GCMG15]. homophily [ABS^{+}12]. Host [HGPS11]. Host-Based [HGPS11]. HOT [LFAH24]. HTML [MDG19, MMH13, RDW^{+}07, UCFL08]. HTTP [PMGO18]. Huge [PBSO14]. Human [ESTD24, MA22, SHHS17]. Humans [GFTC19]. Hybrid [OACU13, ARN12]. Hypergraph [MWHW24]. hypervideo [LV13]. Hypotheses [SHHS17].

IaaS [MFB21]. Identification

[HGPS11, JTLM24, AY10]. Identifying [LCK^{+}12, SSPÇ17, WQG^{+}21]. Identity [BCF16, SKGY14, ZZL^{+}24]. Illness [PWH16]. Image [TC20, UÁM17, ZC23, NKTP13]. Images [UBL^{+}24, Uzu24]. Imaginary [AKJ^{+}18]. Impartial [BCF16]. implementation [CHC13]. Implementing [GTK11]. Implications [JKH^{+}12, WSPZ12, TM09]. Implicit [ZXC^{+}24, DGP09]. Imply [PMGO18]. Improve [SHKK14]. Improving [GCMG15, HHBT20, RSC^{+}21, SFJ^{+}23, VGA13, UT11]. In-depth [SAS24]. In-Real-Life [MJ17]. incentives [JGTF10]. INCEPT [SHLS24]. Incidents [UC22]. Incompatibilities [WAP19]. Incompletely [TBBI18]. Incorporating [ZXC^{+}24]. Incremental [CZ21, MA14]. increments [RDJS07]. index [CKJA13]. Indexing [MA14]. Individual [SXM^{+}16, UT11, ZZM^{+}11]. Inference [BGNV10]. Influence [GCH^{+}21, VVB21, XJZL24, ZS24]. Information [Gae18, GWL^{+}23, JKS23, KL17a, LLW^{+}23, MCFL18, MSP^{+}17, MJ17, PMGO18, RGGG18, SB24, ZWH^{+}24, CSCB07, MMB^{+}12]. Initiative [GCND^{+}15]. Injection [CZZ15]. Inner [AH23]. Innovations [LLA^{+}24]. insecure [YW13]. Insertion [PBSO14]. Insights [GAC^{+}11, RKOD22]. Instantaneous [SLP^{+}19]. Institution [CLU16]. Instructions [ML18]. Integrated [RSC^{+}21]. Integrating [DFJ^{+}12, SZSA15, TDKC15, QBC13]. Integration [DBZ^{+}12, EYH20]. Intelligence [DLMI16]. intensive [DFJ^{+}12]. Intention [OSB22, YZZ^{+}24]. Inter [MHG^{+}20, BKJ13]. Inter-app [MHG^{+}20]. inter-source [BKJ13]. Interaction [LZX^{+}24, Tho14, XJZL24, YZZ^{+}24, WBC08]. Interactions [WSPZ12, JWW^{+}13]. interactive [LV13]. Interconnected [Gae18]. interest [AMND^{+}08, MPB20]. Interface [MvDL12]. interfaces [SWL^{+}13]. International [MA22]. Internet [CGM14, DJBO14, CSCB07, DWS^{+}12, FCBC10, JTVM22, UPS^{+}07, WDD15]. Interpretable [WWN^{+}20]. Introduction [AB08, AI07, DK08, ND08, PYWY23, PYWY24, RS11]. Invariant [SÖ23]. inverted [CKJA13]. Investigating [VVB21]. Investigation [FG18, XLWS17].

Investing [WWW^{+}17]. Investment [AL23]. IRLbot [LLWL09]. Isomorphic [Hog17]. Issue [PYWY23, PYWY24, AB08, DK08, RS11]. issues [ND08]. Item [PCCS23, YZZ^{+}24].

JavaScript [CCH^{+}22, YW13]. Joint [SDSS24]. JSAnalyzer [CCH^{+}22]. Judgments [SXM^{+}16].

Keyword [ZB21]. Keywords [CCL^{+}23]. Keywords-enhanced [CCL^{+}23]. KiN [PAL18]. Know [MNB^{+}24]. Knowledge [ELM16, JWJ^{+}18, LZX^{+}24, MTDF18, PAL18, QWW^{+}24, XJZL24, ZTC11, BLW13, GBF^{+}09, QBC13]. knowledge-aware [QBC13]. known [CCZP24]. Kodi [XVWK23].

Label [BWLK10]. Labeled [PPPS18]. Labelling [Hog17]. Labels [PMS^{+}24]. Language [AKZ20, ABLW19, BHW13, FPG15, ZLL^{+}23, DAA13]. Large [AALG22, CLU16, CSLL18, GFTC19, JKH^{+}12, LLA^{+}24, PGAW23, WPC^{+}22, RDM09, XLL^{+}10]. Large-Scale [CLU16, CSLL18, JKH^{+}12, LLA^{+}24, PGAW23, AALG22, GFTC19, WPC^{+}22, RDM09, XLL^{+}10]. Last [BYJ^{+}21]. Latent [KI23, APV10, JWW^{+}13]. Layout [HY23, MMH13, WAP19, WdSdO23, CM12]. LDAP [KV11]. LDoW [RP17]. Leaders [JKS23]. Leak [ABO^{+}16]. Leak-Aware [ABO^{+}16]. Leaning [Hog17]. Learning [AL24, BGNV10, CZZ15, CTC^{+}15, CCL^{+}23, CRPLM11, HY23, MFB21, QWW^{+}24, Ric08, STZL20, SDG^{+}23, WLP^{+}23, YZD^{+}24, YZZ^{+}24, ZM12]. Learning-based [MFB21]. legal [RDM09]. Leniency [LLW12]. Lens [MZBD20]. Less [PMGO18]. Level [UÁM17, DWXL23, ETT08, SIYL08]. Leveraging [SKGY14, WBC08]. Life [MJ17, PWH16]. Limitations [CCFF11]. Linguistic [BZR^{+}21]. Link [BCD^{+}08, KMSB24, LTS^{+}24, WST11, ZWML14]. Link-Based [ZWML14]. Linkage [ZZL^{+}24]. Linked [RP17, TTHS19]. Linking [GCH^{+}18]. links [Jan07]. List [Dav18c]. Literature [ZLKL19]. Local [CBB18, DDB^{+}14, LMLC24, YL08]. Locally [JWJ^{+}18]. Localness [PAL18]. Location [KKSS17, PAL18, ZZM^{+}11]. Location-Based [KKSS17, PAL18]. locations [ZZM^{+}11]. log [AB08, Coo08, PSBY10]. logs [Ric08, VKY10, XLL^{+}10]. Long [BWLK10, PVS19, Ric08]. Long-term [PVS19, Ric08]. Longitudinal [VDM^{+}18]. Look [GCND^{+}15]. Lossless [MWHW24]. Low [UÁM17]. Low-Level [UÁM17].

Machine [HY23]. Made [UC22]. Maintenance [PBSO14]. Makes [SB24]. Management [AL23, DNPR18, SLL^{+}15, DFJ^{+}12, GKS^{+}08]. Manipulation [DLMI16, Lee15, RDTT24]. map [XLL^{+}10]. Mapping [CGM14]. Market [BSW23]. marketing [LAH07]. Markets [SCW^{+}10]. Mashup [SDC14]. Mashups [CMP15, TKS11]. Masses [BMSV18]. Massive [BMSV18]. Massively [ESTD24]. Matching [LZX^{+}24, ETT08]. Matter [UÁM17]. Maximization [ZS24]. Maximum [AH23]. Measure [GZYS16, LCZ^{+}20, PCCS23, WDD15]. Measurement [CR20, PVS19, YW13]. Measurements [BYG11, AY10]. Measures [KKSS17, Lee15, ZLW^{+}22]. Measuring [JTVM22, MA22, WHS13]. Mechanism [CBB17]. Media [AKJ^{+}18, BSW23, CLB^{+}19b, CZ21, JKWN24, LFAH24, PMS^{+}24, RGGG18, ZJ20, ABS^{+}12]. Mediated [Fra16]. Mediates [KL17a]. Memory [ÁL16]. Mental [ALS23]. Merging [DDB^{+}14]. Message [Whi21]. Meta [QWW^{+}24, WLP^{+}23]. Meta-learning [QWW^{+}24]. Meta-structure [WLP^{+}23]. Metadata [UBL^{+}24]. Method [SHHS17]. Methodology [DBZ^{+}12, GAC^{+}11, CM12]. Methods [LCZ^{+}20, RN09]. Metric [WDD15]. Microblogs [CLR^{+}19]. Millennial [RH19]. Minimization [PVK22]. minimum [RDJS07]. Mining [AZ19, BWLK10, GZC^{+}16, PYWY23, PYWY24, SCS^{+}14, WLT^{+}21, PSBY10]. Mises [WWC^{+}24]. Misinformation [PSP23, DQSZ19, WWW^{+}17]. Misogyny [PAC^{+}21]. Mitigate [RDTT24]. Mitigating [MKH22, SIYL08]. Mobile [CCH^{+}22, CSCB07, DV18, LLA^{+}24, PMGO18, SSK^{+}17, WdSdO23]. Mobility [CPX14]. modal [ZC23]. modalities [SMRM07]. Model [AYJS24, CCL^{+}23, DJBO14, Gae18, KMV15, PMOB11, SGJC20, SW11, SMRM07, SFJ^{+}23, VP11, WWN^{+}20, WML23, ZLL^{+}23, BLW13, CKJA13, CM12, FCBC10, NKTP13]. Model-Based [DJBO14, ZLL^{+}23]. Model-directed [SMRM07]. Model-Driven [VP11, SFJ^{+}23, CM12, FCBC10]. Modeling [MSBB10, MCFL18, PKT17, PVK22, RSC^{+}21, TDKC15, VVB21, WGF24, XZ17, ZHD07, DWS^{+}12, UPS^{+}07, RH19]. Modellus [DWS^{+}12]. Models [ACC08, BBBF14, HHBT20, LCZ^{+}20, RKOD22, WZT^{+}17, WdSdO23, ZM12]. modes [ZCL^{+}10]. Mold [DLMI16]. Money [SCW^{+}10]. Monitoring [KL10]. MOOCs [GWL^{+}23]. MPEG [UÁM17]. MPEG-7 [UÁM17]. Multi [BDM17, LLW^{+}23, PMS^{+}24, SRG^{+}22, ZZW^{+}24, ZC23, ZCC^{+}23]. Multi-behavior [ZZW^{+}24]. Multi-Channel [LLW^{+}23]. Multi-Dimensional [BDM17]. Multi-Labels [PMS^{+}24]. Multi-modal [ZC23]. Multi-Party [SRG^{+}22]. Multi-Task [ZCC^{+}23]. Multidevice [CMP15]. Multimodal [AL24]. Multiplayer [ESTD24]. Multiple [CB20, ZXC^{+}24]. Multiplex [JXCX21].

Multirelational [VBM17]. Multiresolution [LMLC24]. multitier [UPS^{+}07]. MultiView [PBSO14]. MuLX [PMS^{+}24]. MuLX-QA [PMS^{+}24]. Mutable [FGH^{+}16]. Mutexion [MKH22]. Mutually [MKH22]. My [KSG^{+}22]. MyAdChoices [PAAC17]. Myths [SHKK14].

NautiLOD [FPG15]. Naver [PAL18]. Navigating [DWC12, Tho14]. Navigation [MHG^{+}20, RP17, WDD15]. nearest [MKR07]. nearest-neighbor [MKR07]. Neighbor [BCC14, YZZ^{+}24, MKR07]. Network [ALS23, AL24, CLZ^{+}22, GLY^{+}24, JKS23, LTS^{+}24, LZX^{+}24, MWHW24, WLT^{+}21, ZJZ^{+}21, ZLD^{+}21, ZLG^{+}24, ZZW^{+}24, ZCC^{+}23, ZXC^{+}24, YL08]. Networks [AKZ20, ABO^{+}16, AL23, CYL^{+}21, Gae18, GCH^{+}18, GCH^{+}21, GWL^{+}23, JKS23, JXCX21, KL17a, KSG^{+}22, KH15, KMSB24, MSP^{+}17, NRS^{+}22, PSP23, PBSO14, QA14, RUK19, SDSS24, SRG^{+}22, SLL^{+}15, TC20, VBM17, WHGS16, WQG^{+}21, WZJ^{+}24, WSPZ12, WWN^{+}20, ZBG^{+}15, ZS24, ZLW^{+}22, AMND^{+}08, CHC13, Gol09, JWW^{+}13, MMB^{+}12]. Neural [CLZ^{+}22, JKS23, MWHW24, PAC^{+}21, TC20, WWN^{+}20, ZLG^{+}24, ZCC^{+}23, ZXC^{+}24]. News [AL23, RKOD22, SDSS24, DQSZ19, ZLG^{+}24]. Next [LZKN22, SQL^{+}23]. Next-Day [LZKN22]. NFT [SAS24]. Niffler [DWXL23]. NN [XLC20]. Node [GLY^{+}24]. Nodes [Hog17, WQG^{+}21]. Noise [CZZ15]. Non [CCH^{+}22]. Non-critical [CCH^{+}22]. Nonfunctional [FLT15]. Nonlinear [MSP^{+}17]. nonsponsored [Jan07]. Note [YZD^{+}24]. Note-based [YZD^{+}24]. Notices [CLU16]. Novel [GZYS16, LCZ^{+}20, PCCS23]. nuanced [Gol09]. Nucleus [SSPÇ17]. Nudges [RDTT24]. Numbers [AKJ^{+}18].

Obfuscation [RG17]. Objectives [CB20]. Objects [CRPLM11, DDB^{+}14]. obligations [RDM09]. Obscenity [RG17]. Observations [GAC^{+}11]. Obstructing [RG17]. off [FLT15]. Offensive [LFAH24]. Offloading [JJM20]. offs [BYGJ^{+}08]. on-demand [SRRG07]. Online [AKZ20, AYJS24, AKJ^{+}18, AL23, BYJ^{+}21, CZ21, DGS17, ESTD24, Gae18, GCH^{+}18, GCH^{+}21, HGC^{+}18, Hua13, JKS23, JCD^{+}21, KL17a, KSG^{+}22, LLW12, Lee15, MNB^{+}24, MA22, PAAC17, SKGY14, SCB17, SCW^{+}10, Tho14, WH14, WSPZ12, XZ17, ZBG^{+}15, Gol09, JWW^{+}13, JGTF10]. Ontologies [MTDF18]. OP [GTK11]. OP2 [GTK11]. Open [KMV15, LHJL13]. OpenNet [GCND^{+}15]. Operable [BCF16]. Operations [SGJC20]. Opinion [DLMI16, JKS23, UÁM17, XZ17]. Opportunities [DA15]. Optimal [VKY10]. Optimisation [FPWC22]. optimization [YB08]. Optimizing [WYJ^{+}18]. Organizational [GAC^{+}11]. Organizer [ZJZ^{+}21]. oriented [AMH^{+}21, DK08, GLY^{+}24, LMJ10, RCS^{+}08, ZHD07]. OSN [BCF16, RHLC17]. OSN-Based [RHLC17]. OSNs [KI23]. Othering [ABLW19]. Outline [YZD^{+}24]. outsourced [SSL09, SSL09]. Outsourcing [SGJC20]. Overview [CHT^{+}23].

P2P [HGPS11]. P3P [RDM09]. Packages [ZCC^{+}23]. Packing [MMMD16]. Page [BHW13, EYH20, UÁM17, WYJ^{+}18, WST11]. Pages [BKNW23, CCH^{+}22, EYH16, HY23, Uzu24, WHS13, LLWL09, XLH^{+}09]. Pairwise [YMZ19]. PaN [RP17]. Pandemic [CCZP24]. Paradigm [PCCS23]. parallel [WJH13]. parsing [SWL^{+}13]. Part [PYWY23, PYWY24]. partitioning [CKJA13]. Partners [TCM^{+}18]. Party [SRG^{+}22]. Passage [WMTO23]. path [WJH13]. pattern [ZHD07]. PatternRank [XLC20]. Patterns [BAP13, ZHD07]. PeaCE [FGH^{+}16]. PeaCE-Ful [FGH^{+}16]. Penetrating [YMZ19]. Penetrating-rank [YMZ19]. People [AKJ^{+}18, KL17b]. Performance [BCC14, CCZP24, JKH^{+}12, CCFF11, CLB19a]. Periodicity [DGS17]. Personality [KSG^{+}22]. Personalized [FLT15, QRD^{+}22, WML23, ZLG^{+}24]. Personas [AKJ^{+}18]. Perspective [MCFL18, Coo08]. Phases [AMH^{+}21]. Phone [RH19]. Photo [SCB17]. PIDKG [XJZL24]. Piggybacking [CLR^{+}19]. place [RN09]. Platform [KL10, WAP19]. Platforms [BPG^{+}22, PKT17, SDC14, WWW^{+}17, YZD^{+}24]. POI [SQL^{+}23]. Point [MPB20]. Point-of-interest [MPB20]. Polarization [DQSZ19]. Policies [JTVM22, WSL^{+}19, RDM09]. Policy [CRB18, Coo08]. Political [Lee15]. popular [WBC08]. Popularity [HHBT20, SLP^{+}19]. portal [AY10]. Portals [CLB19a]. Post [ZLD^{+}21]. Post-Syncing [ZLD^{+}21]. Posts [PMS^{+}24, SHLS24, WYY^{+}15]. Potential [DQSZ19]. Power [LLW^{+}23]. Practical [ALR^{+}11, LCLQ19]. Practice [ZLD^{+}21]. practices [YW13]. Pre [SQL^{+}23, ZLL^{+}23]. Pre-trained [ZLL^{+}23]. Pre-Training [SQL^{+}23]. Predict [SLP^{+}19, ZJZ^{+}21]. Predictability [ESTD24, HKH^{+}16]. Predicting [CBB18, KSG^{+}22]. Prediction [BCC14, BSW23, GCH^{+}21, HHBT20, HKH^{+}16, KMSB24, LZKN22, LTS^{+}24, TC20, WZZ^{+}16, ZZW^{+}24, ZC23, ABS^{+}12]. Predictive [WML23]. Predictor [CCZP24]. Preference [FLT15]. Prefetching [OACU13]. Presentation [RP17, WYJ^{+}18]. preserving [PSBY10]. PRF [WMTO23]. primitives [ZHD07]. Print [KL17b]. Privacy [BCF16, CLU16, JTVM22, KI23, PSBY10, SCB17, TC20, WSL^{+}19, ZC23, Coo08, RDM09]. Privacy-Aware [BCF16]. privacy-enhancing [Coo08]. Privacy-preserving [PSBY10]. private [CHC13]. Probabilistic [XZ17, ZYZ16, MSBB10]. Problem [MMMD16, AMND^{+}08, CM12]. Process [CJQ^{+}21, RHLC17, LMJ10, VVCD13, ZHD07]. process-driven [ZHD07]. Processes [CJQ^{+}21, TDKC15]. Processing [CAOU12, FGH^{+}16, CKJA13]. producers [DAA13]. Product [AH23, WML23]. profile [Gol09]. Profiles [BCF16, CPX14]. Profiling [PMGO18]. Progress [HR13]. project [LHJL13]. Promise [LFAH24]. promoters [MKR07]. Promoting [ZLZL16]. Propagating [XJZL24, ZWML14]. Propagation [JBWR20, MTDF18, RKOD22, SW11]. Properties [AALG22, WOM22]. proposals [CCFF11]. Protect [CTC^{+}15]. Protecting [JBB^{+}09]. protection [PSBY10]. Protocol [DBZ^{+}12, Fra16]. protocols [RCS^{+}08]. Provenance [KMV15]. Provision [SLL^{+}15]. Proxy [PVS19, RDJS07]. Pseudo [WMTO23]. Pseudo-Relevance [WMTO23]. Public [DLMI16]. publication [PT09]. Publisher [SDSS24]. Pulls [SAS24]. Purchase [LZKN22].

Q2P [WMS^{+}16]. QA [PMS^{+}24]. QoS [ARN12, CB20, CP09, Hua13, WZZ^{+}16, YZL07, ZYZ16]. QoS-aware [CB20]. QoS-based [CP09]. Quality [FG18, LLW12, RDW^{+}16, GKS^{+}08, MSBB10]. Quantitative [WDD15]. Queries [BWLK10, CBB18, CASN13, FPWC22, PPPS18, XLC20, ZB21, GBF^{+}09, Jan07]. Query [BWLK10, CAOU12, DDB^{+}14, HKH^{+}16, OAU11, SXM^{+}16, WBdR12, WST11, WMS^{+}16, AB08, CKJA13, Coo08, PSBY10, Ric08, SWL^{+}13, VKY10, YB08]. Query-Dependent [WST11]. Query-URL [SXM^{+}16]. Queryable [NRS^{+}22]. Querying [TBBI18, QBC13]. Question [SB16]. quite [WOHM08].

Random

[BZA24, JXCX21, KH15, ZBG^{+}15]. rank [YMZ19]. Ranking [BDM17, CZZ15, CSLL18, SHKK14, XLC20, ZM12, ZLL^{+}23]. Ransomware [WPC^{+}22]. Rare [BWLK10]. Rates [SLP^{+}19]. Rating [LLW12]. Ratings [AL24, LLSL18, SCS^{+}14, MKR07, UT11]. Rationale [PMS^{+}24]. Rationales [DA15]. RDF [DNPR18, Hog17, LCLQ19, ZTC11]. RDF/S [ZTC11]. Re [BKNW23, JTLM24]. Re-experience [BKNW23]. Re-Identification [JTLM24]. Readers [PGAW23]. readership [APV10]. Real [AKJ^{+}18, DWXL23, HGPS11, MJ17]. Real-Time [HGPS11, DWXL23]. Reality [SHKK14]. rebinding [JBB^{+}09]. Recommendation [BPG^{+}22, CLZ^{+}22, CCL^{+}23, GWL^{+}23, LCZ^{+}20, LLSL18, MWHW24, MPB20, PSP23, QRD^{+}22, RSC^{+}21, SQL^{+}23, TCM^{+}18, VBM17, WZZ^{+}16, WOM22, WCF^{+}23, WLCG20, WGF24, XJZL24, YZZ^{+}24, ZLG^{+}24, ZCC^{+}23, ZXC^{+}24, LLM13, SZG11, VVCD13]. recommendations [UT11]. Recommender [GZYS16, ZWH^{+}24, CCFF11, RS11, SZG11]. Recommending [ZZM^{+}11]. records [BLW13]. Recovery [BZR^{+}21]. Recurrent [LLA^{+}24]. Reddit [MZBD20]. Reduce [DV18]. Regular [BGNV10]. Reinforced [GWL^{+}23, YYL^{+}23]. Reinforcement [CCL^{+}23, SDG^{+}23]. Relapse [BZR^{+}21]. Relating [SCW^{+}10]. Relational [HY23, KSG^{+}22, SDSS24]. Relationship [WLT^{+}21]. Relevance [SXM^{+}16, WMTO23, BKJ13]. Relevant [Uzu24]. Reliable [SLL^{+}15]. Remotely [KL10]. Rendered [HY23]. Replay [BKNW23]. Reporting [JGTF10]. Repositories [BBBF14]. Representation [ÁL16, STZL20, WCF^{+}23]. Representations [CN10, SHLS24]. Representing [AKJ^{+}18]. Reputation [PKT17, SCW^{+}10, XLWS17]. ReputationPro [ZWZL15]. Requirements [VP11]. Researcher [GCMG15]. Resilient [SLL^{+}15]. Responses [MP23]. Responsive [WdSdO23]. RESTful [BAP13]. Restricted [JXCX21]. Result [NAS16, OAU11, OACU13]. Results [CR20, DDB^{+}14, BKJ13]. Retention [SLP^{+}19]. Retrieval [MCFL18, WMTO23, YMZ19, ZB21, ZLL^{+}23, NKTP13]. Reusable [ELM16]. Reverse [AH23]. Review [JCD^{+}21, PCCS23, RKOD22, WOM22, ZLKL19, JGTF10]. Reviewers [Dav18c, Dav18a]. Reviews [LLSL18, XZ17]. Rich [WDD15, FCBC10, CGM14, DJBO14]. Risk [AL23]. Roaming [MHG^{+}20]. Robust [BLW13, PKT17, WWN^{+}20]. Role [SDSS24]. Role-relational [SDSS24]. Roles [RUK19, MKR07]. RoSGAS [YYL^{+}23]. Rug [SAS24]. Rule [TBBI18, WLCG20]. Rule-Based [TBBI18]. runtime [ACC08].

S [PMGO18, ZTC11]. Sample

[BYJ^{+}21, WCZ15]. Sampled [WCZ15]. Sampling [JXCX21, ZBG^{+}15]. Scalability [WLCG20]. Scalable [AMND^{+}08, CBB17, NAS16, CCFF11]. Scale [CLU16, CSLL18, JKH^{+}12, LLA^{+}24, PGAW23, SÖ23, WSL^{+}19, ZJ20, AALG22, GFTC19, RDM09, WPC^{+}22, XLL^{+}10]. Scale-Invariant [SÖ23]. Scaling [LLWL09]. Scanpath [EYH16]. Scanpaths [EYH16]. Scenarios [JCD^{+}21, PMGO18]. Schemas [BGNV10]. scheme [AADS13]. Scoring [GZC^{+}16, KI23, PPPS18]. Scouts [MKR07]. Scraping [Uzu24]. Screens [MJ17]. Search [AH23, BWLK10, BYG11, BBBF14, CBB18, CZZ15, CAOU12, CR20, CRPLM11, DDB^{+}14, DGS17, HKH^{+}16, LMLC24, MA14, MP23, NAS16, OAU11, OACU13, PWH16, RDTT24, SSL09, SHKK14, TWH14, WYJ^{+}18, WH14, WML23, XLC20, YYL^{+}23, ZM12, ZLL^{+}23, BYGJ^{+}08, CSCB07, GBF^{+}09, LJP^{+}13, ND08, VKY10, WBC08, WH13, XLL^{+}10]. Search-as-a-service [SSL09]. Searching [BDM17, DWC12]. Second [MJ17, OACU13]. section [ND08]. Security [AADP19, CRB18, AADS13]. Selecting [JKWN24]. Selection [BCC14, FLT15, WOM22, WdSdO23, YZL07]. Self [AMH^{+}21, YYL^{+}23, LSC^{+}08]. Self-healing [AMH^{+}21]. self-similarity [LSC^{+}08]. Self-supervised [YYL^{+}23]. Semantic [ELM16, LHJL13, LLM13, LZX^{+}24, WMTO23, ZB21, AMND^{+}08, ETT08, PT09]. Semantics [CRB18, RN09]. Semantics-Based [CRB18]. Semi [QWW^{+}24, BLW13]. semi-structured [BLW13]. Semi-supervised [QWW^{+}24]. Sentiment [CASN13, VCK14, LCZ^{+}20]. Sentiment-Focused [VCK14]. Sequence [SLM13]. Sequential [CLZ^{+}22, MFB21, YZZ^{+}24]. series [VKY10]. Servers [JKH^{+}12, SIYL08]. Service [AMH^{+}21, BAP13, CJQ^{+}21, CBB17, CB20, ELM16, FLT15, STZL20, SLL^{+}15, SZSA15, WZZ^{+}16, WPB13, XLWS17, ZLKL19, ZYZ16, ARN12, CP09, DK08, GKS^{+}08, LMJ10, MPvdA^{+}10, RCS^{+}08, SDN08, SSL09, SIYL08, YB08, ZHD07]. Service-oriented [AMH^{+}21, LMJ10, RCS^{+}08, ZHD07]. Services [AADP19, BCGL17, GAC^{+}11, Hua13, PMOB11, SFJ^{+}23, AADS13, BEP^{+}08, ETT08, GKS^{+}08, KWLA13, LLM13, PT09, QBC13, YZL07]. Session [MWHW24, WGF24, ZWH^{+}24, ZM12, ZCC^{+}23]. Session-Based [WGF24, ZWH^{+}24, ZCC^{+}23, MWHW24]. Session-Context [ZM12]. Set [XLC20]. Sexism [PAC^{+}21]. sGrow [SÖ23]. Shared [KI23, RHS09]. Sharing [MAY^{+}11, MJ17, PAL18, WWN^{+}20, YZD^{+}24]. SHGCN [ZZW^{+}24]. Shifts [PWH16]. Shot [LZX^{+}24]. Should [AST19, WCZ15]. Side [DA15, KL10, MKH22]. Side-Channel [MKH22]. sided [BPG^{+}22]. Signed [CYL^{+}21]. Signs [UC22]. Similar [BYKS09]. similarities [UCFL08]. Similarity [GZYS16, KKSS17, LCZ^{+}20, WZJ^{+}24, YMZ19, Gol09, LSC^{+}08]. Simpler [PMGO18]. Simplifying [CCH^{+}22]. Simulating [MCFL18]. SIP [DBZ^{+}12]. Site [GCH^{+}18, GCH^{+}21, ZZL^{+}24]. Sites [FG18, ZLD^{+}21]. Situational [RGGG18]. Size [KH15]. Slot [AL24]. small [XLH^{+}09]. Smart [DWXL23, KWLA13]. SMINT [WWN^{+}20]. SMS [OSB22]. Snapshot [JJM20]. Snippet [LCLQ19]. SOAP [DBZ^{+}12]. SOC [DK08]. Social [AKZ20, ABO^{+}16, AKJ^{+}18, AL24, BSW23, CLB^{+}19b, CZ21, DZS^{+}16, Gae18, GCH^{+}18, GCH^{+}21, GZC^{+}16, JKS23, JKWN24, KL17a, KSG^{+}22, KH15, LFAH24, MSP^{+}17, NRS^{+}22, PSP23, PMS^{+}24, PBSO14, RUK19, RGGG18, SCS^{+}14, SW11, SKGY14, SRG^{+}22, SDG^{+}23, WQG^{+}21, WSPZ12, YYL^{+}23, ZBG^{+}15, ZS24, ZJZ^{+}21, ZLD^{+}21, ZXC^{+}24, ZJ20, ABS^{+}12, AMND^{+}08, CHC13, Gol09, JWW^{+}13, SZG11]. Socially [ZZW^{+}24]. software [LLM13]. Solution [MMMD16]. source [BKJ13, GBF^{+}09]. Sources [DNPR18, BKJ13]. Spam [CLR^{+}19, LCK^{+}12, OSB22, SSK^{+}17, ZWML14, ZLZL16, BCD^{+}08, KEG^{+}08, UCFL08]. Spans [PMS^{+}24]. SPARQL [FPWC22]. Spatial [WZZ^{+}16]. Spatial-Temporal [WZZ^{+}16]. Spatio [SXY^{+}23]. Spatio-Temporal [SXY^{+}23]. Special [PYWY23, PYWY24, AB08, DK08, ND08, RS11]. Specialized [CRPLM11, RUK19]. Specific [DBZ^{+}12, SDC14]. Specification [VP11, MPvdA^{+}10]. Spectral [LMLC24]. Speech [ABLW19]. splogs [LSC^{+}08]. sponsored [Jan07]. SpotSpam [OSB22]. Spotting [UC22]. Spread [UC22]. Sprite [MMMD16]. Stamped [SLM13]. Standardized [CLU16]. start [GCH^{+}21, MPB20]. State [MvDL12]. Stated [RDM09]. statistical [SWL^{+}13]. Stock [BSW23, CLR^{+}19]. Storage [ACC24, SSL09]. Storytelling [JKWN24]. Strategies [OAU11]. Strategy [CJQ^{+}21]. Stream [WCZ15, ZBG^{+}15]. Streaming [NRS^{+}22, SÖ23, SRRG07]. Strength [SÖ23]. Stress [RUK19]. Structural [QA14, WQG^{+}21, MK12]. Structure [RUK19, BLW13, SMB^{+}07, WLP^{+}23]. structure-knowledge-driven [BLW13]. Structured [BZR^{+}21, YZD^{+}24, BLW13, XLH^{+}09]. Studies [ZLKL19]. Study [BHMW11, BHW13, CR20, CGM14, DZS^{+}16, GZC^{+}16, KL17b, LLA^{+}24, MNB^{+}24, PAL18, PMGO18, RHLC17, CSCB07, RDM09, TM09, WOHM08, XLL^{+}10, YW13]. Studying [ZLD^{+}21]. style [UCFL08]. subcommunities [APV10]. Subgraphs [SSPÇ17]. Subjective [CCZP24, RDW^{+}16]. subpages [XLH^{+}09]. Summarizing [JKWN24, RGGG18]. Summary [JCD^{+}21]. Supervised [CTC^{+}15, JCD^{+}21, QWW^{+}24, YYL^{+}23]. Support [RDTT24]. Supporting [RCS^{+}08, ACC08]. Survey [AKZ20, CLB19a, LBBA20, SB16, VP11, Coo08]. Sybil [FKW^{+}21]. Syncing [ZLD^{+}21]. System [DZS^{+}16, MKH22, PKT17]. Systematic [CGM14]. Systems [AMH^{+}21, GZYS16, JCD^{+}21, LLW12, LLW^{+}23, SB24, WCF^{+}23, ZWH^{+}24, ACC08, CCFF11, KEG^{+}08, RS11, RHS09, SZG11].

Table [MMH13, ZB21]. Tables [HY23]. tag [SZG11]. Tagging [DZS^{+}16, KEG^{+}08, RHS09]. tags [DKM^{+}07, RN09]. Target [ZWML14]. Targets [DQSZ19]. Task [ZCC^{+}23]. Team [ESTD24, RUK19]. Teams [ABO^{+}16]. Techniques [BHW13, FPWC22, CCFF11, Coo08]. Technology [AB08]. Template [AST19]. Templates [WMS^{+}16]. Temporal [LLSL18, SXY^{+}23, WZZ^{+}16, LSC^{+}08]. Temporally [JWJ^{+}18]. Ten [CGM14]. term [CKJA13, PVS19, Ric08]. term-based [CKJA13]. Test [AADP19, AADS13]. Test-Based [AADP19, AADS13]. Testing [BZA24, MZBD20, TM09]. Text [ALS23, MA14, SHLS24, WLP^{+}23, BYKS09]. Textual [BBBF14]. Their [RUK19, MNB^{+}24, WSPZ12]. Theory [PYWY23, PYWY24, QA14]. Things [MCFL18]. Threats [ABLW19]. Threats-based [ABLW19]. Time [HGPS11, LCZ^{+}20, NRS^{+}22, SLM13, WZZ^{+}16, DWXL23, DKM^{+}07, VKY10]. Time-aspect-sentiment [LCZ^{+}20]. Time-aware [WZZ^{+}16]. Time-evolving [NRS^{+}22]. time-series [VKY10]. Time-Stamped [SLM13]. tomorrow [DWC12]. Tool [CCH^{+}22, YZD^{+}24]. Tools [SB24]. Top [PPPS18]. Top-k [PPPS18]. Topic [BHMW11, STZL20, WST11, NKTP13]. Topic-aware [STZL20]. Topics [JTLM24, RDTT24]. Topology [GLY^{+}24]. Tor [BCGL17]. Toxic [LFAH24]. Trackers [FKW^{+}21]. Tracking [ACC24, EYH16, EYH20, UCFL08]. Trade [FLT15, BYGJ^{+}08]. Trade-off [FLT15]. trade-offs [BYGJ^{+}08]. Traffic [PMGO18]. Trails [BWLK10, SHHS17]. trained [ZLL^{+}23]. Training [SQL^{+}23]. Traits [KSG^{+}22]. Transaction [ZWZL15]. Transactions [SZSA15, SDN08, SMRM07]. Transcoding [DA15]. Transducer [TBBI18]. Transfer [QWW^{+}24]. Transformation [BKNW23]. Transformer [WLP^{+}23]. transforming [XLH^{+}09]. Translation [JWJ^{+}18]. Translation-Based [JWJ^{+}18]. Transparency [PAAC17]. transparent [SIYL08]. transportation [ZCL^{+}10]. Travel [CCL^{+}23, ZCC^{+}23]. tree [ÁL16]. Trend [EYH16, SLP^{+}19]. Triangle [GLY^{+}24]. Triangle-oriented [GLY^{+}24]. Triple [ZXC^{+}24]. Trust [Gol09, QA14, SLL^{+}15, ZWML14, ZWZL15, BKJ13, DFJ^{+}12, VVCD13]. trust-based [VVCD13]. Tunnels [MHG^{+}20]. TWEB [Dav18c, Dav18a]. Tweets [KSG^{+}22, UC22]. Twitter [CCZP24, CLR^{+}19, GFTC19, JBWR20, MZBD20, PVK22, RGGG18, WCZ15, ZBG^{+}15, ZLZL16]. Two [AMH^{+}21, BPG^{+}22]. Two-sided [BPG^{+}22]. Type [HGC^{+}18, LLW^{+}23]. Types [AZ19].

U.S. [CLU16]. UI [CMP15]. UI-Centric [CMP15]. Ultimatum [SRG^{+}22]. Uncertain [LYZ23]. Uncovering [CLR^{+}19]. Understand [AYJS24]. Understanding [BZR^{+}21, GCH^{+}18, JWW^{+}13, PSP23, SAS24, SWL^{+}13, ZLD^{+}21, ZCL^{+}10, BYJ^{+}21]. Unified [PT09, WLCG20]. Unknown [MP23]. Unlabeled [GCMG15]. Unsupervised [CSLL18, JCD^{+}21]. Untrustworthy [UBL^{+}24]. Updates [LLA^{+}24]. Upon [JTVM22]. URL [BHMW11, BHW13, MDG19, SXM^{+}16]. URL-Based [BHMW11, BHW13]. URLs [BYKS09]. Usage [SLP^{+}19]. UsageQoS [Hua13]. Use [AST19, FG18, GAC^{+}11, HGPS11, WCZ15, WOHM08]. User [BCC14, BSW23, CBB18, CMP15, CPX14, DZS^{+}16, DGS17, HHBT20, Hua13, LZKN22, MvDL12, PCCS23, PPPS18, PMGO18, SDSS24, UÁM17, WZT^{+}17, WSPZ12, XLC20, YZZ^{+}24, ZZL^{+}24, ZLKL19, DAA13]. User-Based [BCC14]. User-Centric [BSW23]. User-Defined [PPPS18]. User-Item [YZZ^{+}24]. User-Review-Item [PCCS23]. Users [CCZP24, CCC22, DZS^{+}16, GCH^{+}21, MP23, TWH14]. Using [DBZ^{+}12, JKS23, JJM20, ML18, OSB22, RDW^{+}16, Tho14, TC20, UBL^{+}24, UT11, WdSdO23, ZB21, BLW13, GBF^{+}09, LSC^{+}08, LLM13, MSBB10, WAP19, ZHD07]. Utilized [LLW^{+}23]. Utilizing [FKW^{+}21].

validation [MSBB10]. Validity [BCF16]. Value [WWW^{+}17, MK12]. Values [MA22].

Variational [CYL^{+}21, ZCC^{+}23]. verification [MPvdA^{+}10]. Verify [SKGY14]. Versus [SXM^{+}16]. Vertex [PPPS18]. Via [MJ17, CZZ15, JKWN24, KH15, LLW^{+}23, LSC^{+}08, QWW^{+}24, SDSS24, WWC^{+}24, WMS^{+}16]. Video [HHBT20, MAY^{+}11, YZD^{+}24]. Video-sharing [YZD^{+}24]. Videos [HGC^{+}18]. View [PMGO18, ZJZ^{+}21]. Viewers [MJ17]. viewing [VVB21]. Viral [UC22, LAH07]. Virtual [CHC13]. Visual [VVB21, WHS13, DWC12]. Visualization [QRD^{+}22]. Visualizing [DKM^{+}07, JKWN24]. Vital [UC22]. vlHMM [LJP^{+}13]. vocabularies [RHS09]. vs [HY23, RDM09, RDTT24, ZBG^{+}15]. Vulnerability [RDW^{+}07]. Vulnerability-driven [RDW^{+}07].

W [ÁL16]. W-tree [ÁL16]. Walk [JXCX21, KH15]. Warning [DQSZ19]. Watermarking [Fra16]. Weather [ZJZ^{+}21]. Web [AALG22, CCC22, BKJ13, BLW13, DWC12, KWLA13, WJH13, ACC24, AST19, ARN12, ACC08, AL24, AADS13, BHW13, BCD^{+}08, BEP^{+}08, BKNW23, BDM17, BBBF14, CZZ15, CLB19a, CTC^{+}15, CCH^{+}22, CBB17, CB20, CASN13, CHT^{+}23, CN10, CM12, CP09, CMRV10, CSLL18, DFJ^{+}12, DAA13, DA15, DGP09, ETT08, EYH16, EYH20, FTF^{+}18, FPG15, FKW^{+}21, FG18, Fra16, FGH^{+}16, GBF^{+}09, GAC^{+}11, GCND^{+}15, GTK11, HY23, Hua13, Jan07, JJM20, JTVM22, JKWN24, KL10, KL17b, KPED14, LV13, LCK^{+}12, MSBB10, MDG19, MCFL18, MvDL12, MTDF18, MAY^{+}11, MZBD20, NAS16, ND08, NRS^{+}22, OAU11, PMOB11, PYWY23, PYWY24, PT09, QBC13, RDW^{+}16, RDTT24, RP17, RSC^{+}21, RCS^{+}08, SDN08, SMB^{+}07, SB24, SRRG07, STZL20, SCS^{+}14, SIYL08, SMRM07, SFJ^{+}23, TWH14, UÁM17, UCFL08, Uzu24, VP11, VDM^{+}18, VVB21, VCK14]. Web [WZZ^{+}16, WYJ^{+}18, WdSdO23, WPB13, WOHM08, WBC08, WHS13, XLH^{+}09, XLC20, YZL07, YB08, YW13, ZWML14, ZCL^{+}10, ZM12, ZLL^{+}23]. Web-based [ACC08, GAC^{+}11, MAY^{+}11]. Web-Presentation [RP17]. Webgraphs [ÁL16]. Website [HR13]. Weighting [BCC14]. Welcome [Whi21]. Well [CCZP24]. Well-Being [CCZP24]. Who [BYJ^{+}21, RH19]. Whole [WYJ^{+}18]. Whole-Page [WYJ^{+}18]. Wikipedia [DLMI16, PGAW23]. Wild [SSK^{+}17]. Wisdom [LCK^{+}12]. WISeR [BDM17]. within [RDJS07]. without [Uzu24]. Word [BYJ^{+}21, MA22, NKTP13]. workflow [BEP^{+}08]. Workload [JKH^{+}12]. Workloads [MAY^{+}11]. World [LLSL18, Ric08]. Worlds [EYH20]. Worldwide [GCND^{+}15]. WOT [ESTD24]. WWW [SMB^{+}07].

X [CCZP24]. XML [AZ19, BGNV10, CMRV10, MK12].

Yard [KSG^{+}22]. Years [CGM14]. Young [TWH14].

References

Anisetti:2019:TBS

[AADP19] Marco Anisetti, Claudio Ardagna, Ernesto Damiani, and Gianluca Polegri. Test-based security certification of composite services. ACM Transactions on the Web (TWEB), 13(1):3:1–3:??, February 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Anisetti:2013:TBS

[AADS13] Marco Anisetti, Claudio A.

Ardagna, Ernesto Damiani, and Francesco Saonara. A test-based security certification scheme for Web services. ACM Transactions on the Web (TWEB), 7(2):5:1–5:??, May 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Andriamilanto:2022:LSE

[AALG22] Nampoina Andriamilanto, Tristan Allard, Gaëtan Le Guelvouit, and Alexandre Garel. A large-scale empirical analysis of browser fingerprints properties for web authentication. ACM Transactions on the Web (TWEB), 16(1):4:1–4:62, February 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3478026.

Amitay:2008:ISI

[AB08] Einat Amitay and Andrei Broder. Introduction to special issue on query log analysis: Technology and ethics. ACM Transactions on the Web (TWEB), 2(4):18:1–18:??, October 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Alorainy:2019:EAU

[ABLW19] Wafa Alorainy, Pete Burnap, Han Liu, and Matthew L. Williams. “The Enemy Among Us”: Detecting cyber hate speech with threats-based othering language embeddings. ACM Transactions on the Web (TWEB), 13(3):14:1–14:??, October 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3324997.

Amor:2016:DBT

[ABO^{+}16] Iheb Ben Amor, Salima Benbernou, Mourad Ouziri, Zaki Malik, and Brahim Medjahed. Discovering best teams for data leak-aware crowdsourcing in social networks. ACM Transactions on the Web (TWEB), 10(1):2:1–2:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Aiello:2012:FPH

[ABS^{+}12] Luca Maria Aiello, Alain Barrat, Rossano Schifanella, Ciro Cattuto, Benjamin Markines, and Filippo Menczer. Friendship prediction and homophily in social media. ACM Transactions on the Web (TWEB), 6(2):9:1–9:??, May 2012. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Andreolini:2008:MFS

[ACC08] Mauro Andreolini, Sara Casolari, and Michele Colajanni. Models and framework for supporting runtime decisions in Web-based systems. ACM Transactions on the Web (TWEB), 2(3):17:1–17:??, July 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ahmad:2024:EAW

[ACC24] Zubair Ahmad, Samuele Casarin, and Stefano Calzavara. An empirical analysis of Web storage and its applications to Web tracking. ACM Transactions on the Web (TWEB), 18(1):7:1– 7:28, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3623382.

Amagata:2023:RMI

[AH23] Daichi Amagata and Takahiro Hara. Reverse maximum inner product search: Formulation, algorithms, and analysis. ACM Transactions on the Web (TWEB), 17(4):26:1– 26:??, November 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3587215.

Ashman:2007:I

[AI07] Helen Ashman and Arun Iyengar. Introduction. ACM Transactions on the Web (TWEB), 1(1):1:1–1:??, May 2007. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ashman:2012:E

[AIN12] Helen Ashman, Arun Iyengar, and Marc Najork. Editorial. ACM Transactions on the Web (TWEB), 6(2):5:1–5:??, May 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

An:2018:IPR

[AKJ^{+}18] J. An, H. Kwak, S. Jung, J. Salminen, M. Admad, and B. Jansen. Imaginary people representing real numbers: Generating personas from online social media data. ACM Transactions on the Web (TWEB), 12(4):27:1– 27:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Abulaish:2020:SFL

[AKZ20] Muhammad Abulaish, Ashraf Kamal, and Mohammed J. Zaki. A survey of figurative language and its computational detection in online social networks. ACM Transactions on the Web (TWEB), 14(1):3:1– 3:52, February 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3375547.

Avila:2016:WTC

[ÁL16] Bruno T. Ávila and Rafael D.

Lins. W-tree: a compact external memory representation for webgraphs. ACM Transactions on the Web (TWEB), 10(1):6:1– 6:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ang:2023:IRM

[AL23] Gary Ang and Ee-Peng Lim. Investment and risk management with online news and heterogeneous networks. ACM Transactions on the Web (TWEB), 17(2):8:1–8:??, May 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532858.

Ang:2024:LDM

[AL24] Gary Ang and Ee-Peng Lim. Learning dynamic multimodal network slot concepts from the Web for forecasting environmental, social and governance ratings. ACM Transactions on the Web (TWEB), 18(3): 38:1–38:??, August 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3663674.

Alzoubi:2011:PAA

[ALR^{+}11] Hussein A. Alzoubi, Seungjoon Lee, Michael Rabinovich, Oliver Spatscheck, and Jacobus Van Der Merwe. A practical architecture for an Anycast CDN. ACM Transactions on the Web (TWEB), 5(4):17:1–17:??, October 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ahmed:2023:GAN

[ALS23] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Graph attention network for text classification and detection of mental disorder. ACM Transactions on the Web (TWEB), 17(3):19:1–19:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3572406.

Alhosban:2021:TPS

[AMH^{+}21] Amal Alhosban, Zaki Malik, Khayyam Hashmi, Brahim Medjahed, and Hassan Al-Ababneh. A two phases selfhealing framework for serviceoriented systems. ACM Transactions on the Web (TWEB), 15(2):7:1–7:25, May 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450443.

Aleman-Meza:2008:SSA

[AMND^{+}08] Boanerges Aleman-Meza, Meenakshi Nagarajan, Li Ding, Amit Sheth, I. Budak Arpinar, Anupam Joshi, and Tim Finin. Scalable semantic analytics on social networks for addressing the problem of conflict of interest detection. ACM Transactions on the Web (TWEB), 2(1):7:1– 7:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Anonymous:2015:E

[Ano15] Anonymous. Editorial. ACM Transactions on the Web (TWEB), 9(2):6:1–6:??, May 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Adams:2010:DLS

[APV10] Brett Adams, Dinh Phung, and Svetha Venkatesh. Discovery of latent subcommunities in a blog’s readership. ACM Transactions on the Web (TWEB), 4 (3):12:1–12:??, July 2010. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Alrifai:2012:HAE

[ARN12] Mohammad Alrifai, Thomas Risse, and Wolfgang Nejdl. A hybrid approach for efficient Web service composition with end-to-end QoS constraints. ACM Transactions on the Web (TWEB), 6(2):7:1–7:??, May 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Alarte:2019:WWT

[AST19] Julián Alarte, Josep Silva, and Salvador Tamarit. What Web template extractor should I use? A benchmarking and comparison for five template extractors. ACM Transactions on the Web (TWEB), 13(2):9:1–9:??, April 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3316810.

Almishari:2010:APD

[AY10] Mishari Almishari and Xiaowei Yang. Ads-portal domains: Identification and measurements. ACM Transactions on the Web (TWEB), 4(2):4:1– 4:??, April 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Agarwal:2024:GBC

[AYJS24] Vibhor Agarwal, Anthony P.

Young, Sagar Joglekar, and Nishanth Sastry. A graphbased context-aware model to understand online conversations. ACM Transactions on the Web (TWEB), 18(1):10:1– 10:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3624579.

Athanasopoulos:2019:MAX

[AZ19] Dionysis Athanasopoulos and Apostolos Zarras. Mining abstract XML data-types. ACM Transactions on the Web (TWEB), 13(1):2:1–2:??, February 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bellido:2013:CFP

[BAP13] Jesus Bellido, Rosa Alarcón, and Cesare Pautasso. Controlflow patterns for decentralized RESTful service composition. ACM Transactions on the Web (TWEB), 8(1):5:1–5:??, December 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bislimovska:2014:TCB

[BBBF14] Bojana Bislimovska, Alessandro Bozzon, Marco Brambilla, and Piero Fraternali. Textual and content-based search in repositories of Web application models. ACM Transactions on the Web (TWEB), 8(2):11:1–11:??, March 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bellogin:2014:NSW

[BCC14] Alejandro Bellogín, Pablo Castells, and Iván Cantador. Neighbor selection and weighting in user-based collaborative filtering: a performance prediction approach. ACM Transactions on the Web (TWEB), 8(2):12:1–12:??, March 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Becchetti:2008:LAW

[BCD^{+}08] Luca Becchetti, Carlos Castillo, Debora Donato, Ricardo Baeza-Yates, and Stefano Leonardi. Link analysis for Web spam detection. ACM Transactions on the Web (TWEB), 2(1):2:1– 2:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bahri:2016:CCO

[BCF16] Leila Bahri, Barbara Carminati, and Elena Ferrari. COIPcontinuous, operable, impartial, and privacy-aware identity validity estimation for OSN profiles. ACM Transactions on the Web (TWEB), 10(4):23:1– 23:??, December 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bernaschi:2017:EAT

[BCGL17] Massimo Bernaschi, Alessandro Celestini, Stefano Guarino, and Flavio Lombardi. Exploring and analyzing the Tor hidden services graph. ACM Transactions on the Web (TWEB), 11 (4):24:1–24:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bianchini:2017:WMD

[BDM17] Devis Bianchini, Valeria De Antonellis, and Michele Melchiori. WISeR: a multidimensional framework for searching and ranking Web APIs. ACM Transactions on the Web (TWEB), 11(3):19:1– 19:??, July 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Belhajjame:2008:AAW

[BEP^{+}08] Khalid Belhajjame, Suzanne M.

Embury, Norman W. Paton, Robert Stevens, and Carole A. Goble. Automatic annotation of Web services based on workflow definitions. ACM Transactions on the Web (TWEB), 2(2):11:1– 11:??, April 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bex:2010:LDR

[BGNV10] Geert Jan Bex, Wouter Gelade, Frank Neven, and Stijn Vansummeren. Learning deterministic regular expressions for the inference of schemas from XML data. ACM Transactions on the Web (TWEB), 4(4):14:1–14:??, September 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Baykan:2011:CSF

[BHMW11] Eda Baykan, Monika Henzinger, Ludmila Marian, and Ingmar Weber. A comprehensive study of features and algorithms for URL-based topic classification. ACM Transactions on the Web (TWEB), 5(3):15:1–15:??, July 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Baykan:2013:CST

[BHW13] Eda Baykan, Monika Henzinger, and Ingmar Weber. A comprehensive study of techniques for URL-based Web page language classification. ACM Transactions on the Web (TWEB), 7 (1):3:1–3:??, March 2013. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Balakrishnan:2013:ART

[BKJ13] Raju Balakrishnan, Subbarao Kambhampati, and Manishkumar Jha. Assessing relevance and trust of the deep web sources and results based on inter-source agreement. ACM Transactions on the Web (TWEB), 7(2):11:1–11:??, May 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Berlin:2023:REW

[BKNW23] John Berlin, Mat Kelly, Michael L. Nelson, and Michele C. Weigle. To re-experience the Web: a framework for the transformation and replay of archived Web pages. ACM Transactions on the Web (TWEB), 17(4):28:1– 28:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3589206.

Bing:2013:RDS

[BLW13] Lidong Bing, Wai Lam, and Tak-Lam Wong. Robust detection of semi-structured web records using a DOM structureknowledge-driven model. ACM Transactions on the Web (TWEB), 7(4):21:1–21:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Boldi:2018:BMC

[BMSV18] Paolo Boldi, Andrea Marino, Massimo Santini, and Sebastiano Vigna. BUbiNG: Massive crawling for the masses. ACM Transactions on the Web (TWEB), 12(2):12:1– 12:26, June 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/citation.cfm?
doid=3176641.3160017.

Biswas:2022:TFR

[BPG^{+}22] Arpita Biswas, Gourab K.

Patro, Niloy Ganguly, Krishna P. Gummadi, and Abhijnan Chakraborty. Toward fair recommendation in two-sided platforms. ACM Transactions on the Web (TWEB), 16(2): 8:1–8:34, May 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3503624.

Bouadjenek:2023:UCA

[BSW23] Mohamed Reda Bouadjenek, Scott Sanner, and Ga Wu. A user-centric analysis of social media for stock market prediction. ACM Transactions on the Web (TWEB), 17(2): 9:1–9:??, May 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3532856.

Bailey:2010:MHQ

[BWLK10] Peter Bailey, Ryen W. White, Han Liu, and Giridhar Kumaran. Mining historic query trails to label long and rare search engine queries. ACM Transactions on the Web (TWEB), 4(4):15:1–15:??, September 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Bar-Yossef:2011:ESE

[BYG11] Ziv Bar-Yossef and Maxim Gurevich. Efficient search engine measurements. ACM Transactions on the Web (TWEB), 5(4):18:1–18:??, October 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Baeza-Yates:2008:DTO

[BYGJ^{+}08] Ricardo Baeza-Yates, Aristides Gionis, Flavio P. Junqueira, Vanessa Murdock, Vassilis Plachouras, and Fabrizio Silvestri.

Design trade-offs for search engine caching. ACM Transactions on the Web (TWEB), 2(4):20:1–20:??, October 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Boschi:2021:WLW

[BYJ^{+}21] Gioia Boschi, Anthony P.

Young, Sagar Joglekar, Chiara Cammarota, and Nishanth Sastry. Who has the last word? Understanding how to sample online discussions. ACM Transactions on the Web (TWEB), 15(3):12:1–12:25, July 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3452936.

Bar-Yossef:2009:DCD

[BYKS09] Ziv Bar-Yossef, Idit Keidar, and Uri Schonfeld. Do not crawl in the DUST: Different URLs with Similar Text. ACM Transactions on the Web (TWEB), 3 (1):3:1–3:??, January 2009. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Belhadi:2024:RTE

[BZA24] Asma Belhadi, Man Zhang, and Andrea Arcuri. Random testing and evolutionary testing for fuzzing GraphQL APIs. ACM Transactions on the Web (TWEB), 18(1):14:1– 14:??, February 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3609427.

Bailey:2021:SLA

[BZR^{+}21] Shawn Bailey, Yue Zhang, Arti Ramesh, Jennifer Golbeck, and Lise Getoor. A structured and linguistic approach to understanding recovery and relapse in AA. ACM Transactions on the Web (TWEB), 15(1):5:1–5:35, January 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3423208.

Cambazoglu:2012:CBQ

[CAOU12] B. Barla Cambazoglu, Ismail Sengor Altingovde, Rifat Ozcan, and Özgür Ulusoy. Cache-based query processing for search engines. ACM Transactions on the Web (TWEB), 6 (4):14:1–14:??, November 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chelaru:2013:ADE

[CASN13] Sergiu Chelaru, Ismail Sengor Altingovde, Stefan Siersdorfer, and Wolfgang Nejdl. Analyzing, detecting, and exploiting sentiment in Web queries. ACM Transactions on the Web (TWEB), 8(1):6:1–6:??, December 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chattopadhyay:2020:QAA

[CB20] Soumi Chattopadhyay and Ansuman Banerjee. QoS-aware automatic Web service composition with multiple objectives. ACM Transactions on the Web (TWEB), 14(3): 12:1–12:38, July 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3389147.

Chattopadhyay:2017:FSM

[CBB17] Soumi Chattopadhyay, Ansuman Banerjee, and Nilanjan Banerjee. A fast and scalable mechanism for Web service composition. ACM Transactions on the Web (TWEB), 11 (4):26:1–26:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Cacheda:2018:CPU

[CBB18] Fidel Cacheda, Roi Blanco, and Nicola Barbieri. Characterizing and predicting users’ behavior on local search queries. ACM Transactions on the Web (TWEB), 12(2):11:1– 11:??, June 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Crichton:2022:HDH

[CCC22] Kyle Crichton, Nicolas Christin, and Lorrie Faith Cranor. How do home computer users browse the web? ACM Transactions on the Web (TWEB), 16(1):3:1–3:27, February 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3473343.

Cacheda:2011:CCF

[CCFF11] Fidel Cacheda, Víctor Carneiro, Diego Fernández, and Vreixo Formoso. Comparison of collaborative filtering algorithms: Limitations of current techniques and proposals for scalable, high-performance recommender systems. ACM Transactions on the Web (TWEB), 5(1):2:1–2:??, February 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chaqfeh:2022:JWD

[CCH^{+}22] Moumena Chaqfeh, Russell Coke, Jacinta Hu, Waleed Hashmi, Lakshmi Subramanian, Talal Rahwan, and Yasir Zaki. JSAnalyzer: a Web developer tool for simplifying mobile Web pages through non-critical JavaScript elimination. ACM Transactions on the Web (TWEB), 16(4):17:1– 17:??, November 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3550358.

Chen:2023:KED

[CCL^{+}23] Lei Chen, Jie Cao, Weichao Liang, Jia Wu, and Qiaolin Ye. Keywords-enhanced deep reinforcement learning model for travel recommendation. ACM Transactions on the Web (TWEB), 17(1):5:1–5:??, February 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570959.

Chen:2024:BPX

[CCZP24] Ninghan Chen, Xihui Chen, Zhiqiang Zhong, and Jun Pang. Bridging performance of X (formerly known as Twitter) users: a predictor of subjective well-being during the pandemic. ACM Transactions on the Web (TWEB), 18(1):15:1– 15:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3635033.

Casteleyn:2014:TYR

[CGM14] Sven Casteleyn, Irene Garrigós, and Jose-Norberto Mazón. Ten years of Rich Internet Applications: a systematic mapping study, and beyond. ACM Transactions on the Web (TWEB), 8 (3):18:1–18:??, June 2014. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Conti:2013:VPS

[CHC13] Mauro Conti, Arbnor Hasani, and Bruno Crispo. Virtual private social networks and a Facebook implementation. ACM Transactions on the Web (TWEB), 7(3):14:1–14:??, September 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chen:2023:FWO

[CHT^{+}23] Chung-Chi Chen, Hen-Hsen Huang, Hiroya Takamura, Makoto P. Kato, and Yu-Lieh Huang. FinTech on the Web: an overview. ACM Transactions on the Web (TWEB), 17(2): 7:1–7:??, May 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3572404.

Cao:2021:CCB

[CJQ^{+}21] Jian Cao, Tingjie Jia, Shiyou Qian, Haiyan Zhao, and Jie Wang. CBPCS: a cache-block-based service process caching strategy to accelerate the execution of service processes. ACM Transactions on the Web (TWEB), 15(1):1:1–1:29, January 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3411494.

Cambazoglu:2013:TBI

[CKJA13] B. Barla Cambazoglu, Enver Kayaaslan, Simon Jonassen, and Cevdet Aykanat. A term-based inverted index partitioning model for efficient distributed query processing. ACM Transactions on the Web (TWEB), 7(3):15:1–15:??, September 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Calegari:2019:WPH

[CLB19a] Patrice Calegari, Marc Levrier, and Paweł Balczyński. Web portals for high-performance computing: a survey. ACM Transactions on the Web (TWEB), 13(1):5:1–5:??, February 2019.

CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chatzakou:2019:DCC

[CLB^{+}19b] Despoina Chatzakou, Ilias Leontiadis, Jeremy Blackburn, Emiliano De Cristofaro, Gianluca Stringhini, Athena Vakali, and Nicolas Kourtellis. Detecting cyberbullying and cyberaggression in social media. ACM Transactions on the Web (TWEB), 13(3):17:1–17:??, October 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3343484.

Cresci:2019:CPU

[CLR^{+}19] Stefano Cresci, Fabrizio Lillo, Daniele Regoli, Serena Tardelli, and Maurizio Tesconi. Cashtag piggybacking: Uncovering spam and bot activity in stock microblogs on Twitter. ACM Transactions on the Web (TWEB), 13(2):11:1– 11:??, April 2019. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3313184.

Cranor:2016:LSE

[CLU16] Lorrie Faith Cranor, Pedro Giovanni Leon, and Blase Ur. A large-scale evaluation of U.S. financial institutions’ standardized privacy notices. ACM Transactions on the Web (TWEB), 10(3):17:1–17:??, August 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chen:2022:DAC

[CLZ^{+}22] Qi Chen, Guohui Li, Quan Zhou, Si Shi, and Deqing Zou. Double attention convolutional neural network for sequential recommendation. ACM Transactions on the Web (TWEB), 16 (4):19:1–19:??, November 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3555350.

Comai:2012:MDM

[CM12] Sara Comai and Davide Mazza. A model-driven methodology to the content layout problem in Web applications. ACM Transactions on the Web (TWEB), 6 (3):10:1–10:38, September 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Cappiello:2015:UCA

[CMP15] Cinzia Cappiello, Maristella Matera, and Matteo Picozzi. A UI-centric approach for the enduser development of multidevice mashups. ACM Transactions on the Web (TWEB), 9(3):11:1– 11:??, June 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Consens:2010:EXW

[CMRV10] Mariano P. Consens, Renée J. Miller, Flavio Rizzolo, and Alejandro A. Vaisman. Exploring XML Web collections with DescribeX. ACM Transactions on the Web (TWEB), 4(3):11:1– 11:??, July 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Claude:2010:FCW

[CN10] Francisco Claude and Gonzalo Navarro. Fast and compact Web graph representations. ACM Transactions on the Web (TWEB), 4(4):16:1–16:??, September 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Cooper:2008:SQL

[Coo08]  Alissa Cooper.  A survey

of query log privacy-enhancing techniques from a policy perspective. ACM Transactions on the Web (TWEB), 2(4):19:1– 19:??, October 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Comuzzi:2009:FQB

[CP09] Marco Comuzzi and Barbara Pernici. A framework for QoSbased Web service contracting. ACM Transactions on the Web (TWEB), 3(3):10:1–10:??, June 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chen:2014:CCU

[CPX14] Xihui Chen, Jun Pang, and Ran Xue. Constructing and comparing user mobility profiles. ACM Transactions on the Web (TWEB), 8(4):21:1–21:??, October 2014. CODEN ????

ISSN 1559-1131 (print), 1559-114X (electronic).

Carpineto:2020:ESA

[CR20] Claudio Carpineto and Giovanni Romano. An experimental study of automatic detection and measurement of counterfeit in brand search results. ACM Transactions on the Web (TWEB), 14(2): 6:1–6:35, April 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3378443.

Calzavara:2018:SBA

[CRB18] Stefano Calzavara, Alvise Rabitti, and Michele Bugliesi. Semantics-based analysis of content security policy deployment. ACM Transactions on the Web (TWEB), 12(2):10:1– 10:??, June 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Curlango-Rosas:2011:SSA

[CRPLM11] Cecilia Curlango-Rosas, Gregorio A. Ponce, and Gabriel A. Lopez-Morteo. A specialized search assistant for learning objects. ACM Transactions on the Web (TWEB), 5(4):21:1–21:??, October 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Church:2007:MIA

[CSCB07] Karen Church, Barry Smyth, Paul Cotter, and Keith Bradley. Mobile information access: a study of emerging search behavior on the mobile Internet. ACM Transactions on the Web (TWEB), 1(1):4:1–4:??, May 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Cui:2018:UDR

[CSLL18] Yi Cui, Clint Sparkman, Hsin-Tsang Lee, and Dmitri Loguinov. Unsupervised domain ranking in large-scale Web crawls. ACM Transactions on the Web (TWEB), 12(4):26:1– 26:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Calzavara:2015:SLA

[CTC^{+}15] Stefano Calzavara, Gabriele Tolomei, Andrea Casini, Michele Bugliesi, and Salvatore Orlando. A supervised learning approach to protect client authentication on the Web. ACM Transactions on the Web (TWEB), 9(3):15:1–15:??, June 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chen:2021:DVE

[CYL^{+}21] Xu Chen, Jiangchao Yao, Maosen Li, Ya Zhang, and Yanfeng Wang. Decoupled variational embedding for signed directed networks. ACM Transactions on the Web (TWEB), 15(1):3:1–3:31, January 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3408298.

Chelmis:2021:DIC

[CZ21] Charalampos Chelmis and Daphney-Stavroula Zois. Dynamic, incremental, and continuous detection of cyberbullying in online social media. ACM Transactions on the Web (TWEB), 15(3):14:1– 14:33, July 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3448014.

Cai:2015:ALW

[CZZ15] Wenbin Cai, Muhan Zhang, and Ya Zhang. Active learning for Web search ranking via noise injection. ACM Transactions on the Web (TWEB), 9(1):3:1– 3:??, January 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Diaz:2015:AWR

[DA15] Oscar Díaz and Cristóbal Arellano. The augmented Web: Rationales, opportunities, and challenges on browser-side transcoding. ACM Transactions on the Web (TWEB), 9(2):8:1– 8:??, May 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Diaz:2013:LEU

[DAA13] Oscar Díaz, Cristóbal Arellano, and Maider Azanza. A language for end-user Web augmentation:

Caring for producers and consumers alike. ACM Transactions on the Web (TWEB), 7 (2):9:1–9:??, May 2013. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Davison:2018:TR

[Dav18a] Brian D. Davison. 2017 TWEB reviewers. ACM Transactions on the Web (TWEB), 12(2): 14:1–14:??, June 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Davison:2018:E

[Dav18b] Brian D. Davison. Editorial.

ACM Transactions on the Web (TWEB), 12(2):8:1–8:??, June 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Davison:2018:LTR

[Dav18c] Brian D. Davison. List of 2016 TWEB reviewers. ACM Transactions on the Web (TWEB), 12(1):7:1–7:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Delac:2012:MSS

[DBZ^{+}12] Goran Delac, Ivan Budiselic, Ivan Zuzak, Ivan Skuliber, and Tomislav Stefanec. A methodology for SIP and SOAP integration using applicationspecific protocol conversion. ACM Transactions on the Web (TWEB), 6(4):15:1–15:??, November 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dragut:2014:MQR

[DDB^{+}14] Eduard C. Dragut, Bhaskar Dasgupta, Brian P. Beirne, Ali Neyestani, Badr Atassi, Clement Yu, and Weiyi Meng. Merging query results from local search engines for georeferenced objects. ACM Transactions on the Web (TWEB), 8(4):20:1–20:??, October 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

DeCapitaniDiVimercati:2012:ITM

[DFJ^{+}12] Sabrina De Capitani Di Vimercati, Sara Foresti, Sushil Jajodia, Stefano Paraboschi, Giuseppe Psaila, and Pierangela Samarati. Integrating trust management and access control in data-intensive Web applications. ACM Transactions on the Web (TWEB), 6(2):6:1– 6:??, May 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dourisboure:2009:ECD

[DGP09] Yon Dourisboure, Filippo Geraci, and Marco Pellegrini. Extraction and classification of dense implicit communities in the Web graph. ACM Transactions on the Web (TWEB), 3(2):7:1– 7:??, April 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Drutsa:2017:PUE

[DGS17] Alexey Drutsa, Gleb Gusev, and Pavel Serdyukov. Periodicity in user engagement with a search engine and its application to online controlled experiments. ACM Transactions on the Web (TWEB), 11(2):9:1–9:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dincturk:2014:MBA

[DJBO14] Mustafa Emre Dincturk, Guy-Vincent Jourdan, Gregor V. Bochmann, and Iosif Viorel Onut. A model-based approach for crawling Rich Internet Applications. ACM Transactions on the Web (TWEB), 8(3):19:1– 19:??, June 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dustdar:2008:ISI

[DK08] Schahram Dustdar and Bernd J.

Krämer. Introduction to special issue on service oriented computing (SOC). ACM Transactions on the Web (TWEB), 2 (2):10:1–10:??, April 2008. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dubinko:2007:VTT

[DKM^{+}07] Micah Dubinko, Ravi Kumar, Joseph Magnani, Jasmine Novak, Prabhakar Raghavan, and Andrew Tomkins. Visualizing tags over time. ACM Transactions on the Web (TWEB), 1 (2):7:1–7:??, August 2007. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Das:2016:MAA

[DLMI16] Sanmay Das, Allen Lavoie, and Malik Magdon-Ismail. Manipulation among the arbiters of collective intelligence: How Wikipedia administrators mold public opinion. ACM Transactions on the Web (TWEB), 10 (4):24:1–24:??, December 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Darari:2018:CMR

[DNPR18] Fariz Darari, Werner Nutt, Giuseppe Pirrò, and Simon Razniewski. Completeness management for RDF data sources. ACM Transactions on the Web (TWEB), 12(3):18:1– 18:??, July 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Vicario:2019:PFN

[DQSZ19] Michela Del Vicario, Walter Quattrociocchi, Antonio Scala, and Fabiana Zollo. Polarization and fake news: Early warning of potential misinformation targets. ACM Transactions on the Web (TWEB), 13(2):10:1–10:??, April 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3316809.

Dutta:2018:CRM

[DV18] Kaushik Dutta and Debra Vandermeer. Caching to reduce mobile app energy consumption. ACM Transactions on the Web (TWEB), 12(1):5:1–5:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Dork:2012:NTW

[DWC12] Marian Dörk, Carey Williamson, and Sheelagh Carpendale. Navigating tomorrow’s web: From searching and browsing to visual exploration. ACM Transactions on the Web (TWEB), 6 (3):13:1–13:??, September 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Desnoyers:2012:MAM

[DWS^{+}12] Peter Desnoyers, Timothy Wood, Prashant Shenoy, Rahul Singh, Sangameshwar Patil, and Harrick Vin. Modellus: Automated modeling of complex Internet data center applications. ACM Transactions on the Web (TWEB), 6(2):8:1–8:??, May 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Du:2023:NRT

[DWXL23] Haohua Du, Yue Wang, Xiaoya Xu, and Mingsheng Liu. Niffler: Real-time device-level anomalies detection in smart home. ACM Transactions on the Web (TWEB), 17(3):16:1– 16:??, August 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3586073.

Doerfel:2016:WUA

[DZS^{+}16] Stephan Doerfel, Daniel Zoller, Philipp Singer, Thomas Niebler, Andreas Hotho, and Markus Strohmaier. What users actually do in a social tagging system: a study of user behavior in BibSonomy. ACM Transactions on the Web (TWEB), 10 (2):14:1–14:??, May 2016. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Eshuis:2016:FCE

[ELM16] Rik Eshuis, Freddy Lécué, and Nikolay Mehandjiev. Flexible construction of executable service compositions from reusable semantic knowledge. ACM Transactions on the Web (TWEB), 10(1):5:1–5:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Emmert-Streib:2024:HTB

[ESTD24] Frank Emmert-Streib, Shailesh Tripathi, and Matthias Dehmer. Human team behavior and predictability in the massively multiplayer online game WOT blitz. ACM Transactions on the Web (TWEB), 18(1):5:1– 5:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3617509.

Elgedawy:2008:CAH

[ETT08] Islam Elgedawy, Zahir Tari, and James A. Thom. Correctnessaware high-level functional matching approaches for semantic Web services. ACM Transactions on the Web (TWEB), 2 (2):12:1–12:??, April 2008. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Eraslan:2016:STA

[EYH16] Sukru Eraslan, Yeliz Yesilada, and Simon Harper. Scanpath trend analysis on Web pages: Clustering eye tracking scanpaths. ACM Transactions on the Web (TWEB), 10(4):20:1– 20:??, December 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Eraslan:2020:BBW

[EYH20] Sukru Eraslan, Yeliz Yesilada, and Simon Harper. “The Best of Both Worlds!”: Integration of Web page and eye tracking data driven approaches for automatic AOI detection. ACM Transactions on the Web (TWEB), 14(1):1:1–1:31, February 2020. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372497.

Fraternali:2010:ERI

[FCBC10] Piero Fraternali, Sara Comai, Alessandro Bozzon, and Giovanni Toffetti Carughi. Engineering rich Internet applications with a model-driven approach. ACM Transactions on the Web (TWEB), 4(2):7:1– 7:??, April 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Fogli:2018:EQU

[FG18] Daniela Fogli and Giovanni Guida. Evaluating quality in use of corporate Web sites: an empirical investigation. ACM Transactions on the Web (TWEB), 12(3):15:1– 15:??, July 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Furche:2016:PFW

[FGH^{+}16] Tim Furche, Giovanni Grasso, Michael Huemer, Christian Schallhart, and Michael Schrefl. PeaCE-Ful Web event extraction and processing as bitemporal mutable events. ACM Transactions on the Web (TWEB), 10(3):16:1–16:??, August 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Flores:2021:UWT

[FKW^{+}21] Marcel Flores, Andrew Kahn, Marc Warrior, Alan Mislove, and Aleksandar Kuzmanovic. Utilizing Web trackers for Sybil defense. ACM Transactions on the Web (TWEB), 15(2): 8:1–8:19, May 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3450444.

Fletcher:2015:EPN

[FLT15] Kenneth K. Fletcher, Xiaoqing F. Liu, and Mingdong Tang. Elastic personalized nonfunctional attribute preference and trade-off based service selection. ACM Transactions on the Web (TWEB), 9(1):1:1– 1:??, January 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Fionda:2015:NFL

[FPG15] Valeria Fionda, Giuseppe Pirrò, and Claudio Gutierrez. NautiLOD: a formal language for the Web of Data graph. ACM Transactions on the Web (TWEB), 9(1):5:1–5:??, January 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Frosini:2022:OTF

[FPWC22] Riccardo Frosini, Alexandra Poulovassilis, Peter T. Wood, and Andrea Calí. Optimisation techniques for flexible SPARQL queries. ACM Transactions on the Web (TWEB), 16(4):16:1– 16:??, November 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3532855.

Frattolillo:2016:BFM

[Fra16] Franco Frattolillo. A buyerfriendly and mediated watermarking protocol for Web context. ACM Transactions on the Web (TWEB), 10(2):9:1– 9:??, May 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Fanou:2018:EAA

[FTF^{+}18] Rodérick Fanou, Gareth Tyson, Eder Leao Fernandes, Pierre Francois, Francisco Valera, and Arjuna Sathiaseelan. Exploring and analysing the African Web ecosystem. ACM Transactions on the Web (TWEB), 12 (4):22:1–22:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gill:2011:COU

[GAC^{+}11] Phillipa Gill, Martin Arlitt, Niklas Carlsson, Anirban Mahanti, and Carey Williamson. Characterizing organizational use of Web-based services: Methodology, challenges, observations, and insights. ACM Transactions on the Web (TWEB), 5(4):19:1–19:??, October 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gaeta:2018:MID

[Gae18]  Rossano Gaeta.  A model

of information diffusion in interconnected online social networks. ACM Transactions on the Web (TWEB), 12(2):13:1– 13:??, June 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gabrilovich:2009:CSQ

[GBF^{+}09] Evgeniy Gabrilovich, Andrei Broder, Marcus Fontoura, Amruta Joshi, Vanja Josifovski, Lance Riedel, and Tong Zhang. Classifying search queries using the Web as a source of knowledge. ACM Transactions on the Web (TWEB), 3(2):5:1– 5:??, April 2009. CODEN ????

ISSN 1559-1131 (print), 1559-114X (electronic).

Gong:2018:UCS

[GCH^{+}18] Qingyuan Gong, Yang Chen, Jiyao Hu, Qiang Cao, Pan Hui, and Xin Wang. Understanding cross-site linking in online social networks. ACM Transactions on the Web (TWEB), 12 (4):25:1–25:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gong:2021:CSP

[GCH^{+}21] Qingyuan Gong, Yang Chen, Xinlei He, Yu Xiao, Pan Hui, Xin Wang, and Xiaoming Fu. Cross-site prediction on social influence for cold-start users in online social networks. ACM Transactions on the Web (TWEB), 15(2):6:1–6:23, May 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3409108.

Gollapalli:2015:IRH

[GCMG15] Sujatha Das Gollapalli, Cornelia Caragea, Prasenjit Mitra, and C. Lee Giles. Improving researcher homepage classification with unlabeled data. ACM Transactions on the Web (TWEB), 9(4):17:1–17:??, October 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gill:2015:CWC

[GCND^{+}15] Phillipa Gill, Masashi Crete-Nishihata, Jakub Dalek, Sharon Goldberg, Adam Senft, and Greg Wiseman. Characterizing Web censorship worldwide: Another look at the OpenNet initiative data. ACM Transactions on the Web (TWEB), 9(1):4:1– 4:??, January 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gilani:2019:LSB

[GFTC19] Zafar Gilani, Reza Farahbakhsh, Gareth Tyson, and Jon Crowcroft. A largescale behavioural analysis of bots and humans on Twitter. ACM Transactions on the Web (TWEB), 13(1):7:1–7:??, February 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gmach:2008:AQS

[GKS^{+}08] Daniel Gmach, Stefan Krompass, Andreas Scholz, Martin Wimmer, and Alfons Kemper. Adaptive quality of service management for enterprise services. ACM Transactions on the Web (TWEB), 2(1):8:1–8:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gao:2024:TOC

[GLY^{+}24] Guangliang Gao, Weichao Liang, Ming Yuan, Hanwei Qian, Qun Wang, and Jie Cao. Triangle-oriented community detection considering node features and network topology. ACM Transactions on the Web (TWEB), 18(1):11:1– 11:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3626190.

Golbeck:2009:TNP

[Gol09] Jennifer Golbeck. Trust and nuanced profile similarity in online social networks. ACM Transactions on the Web (TWEB), 3 (4):12:1–12:??, September 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Grier:2011:DIO

[GTK11] Chris Grier, Shuo Tang, and Samuel T. King. Designing and implementing the OP and OP2 Web browsers. ACM Transactions on the Web (TWEB), 5 (2):11:1–11:??, May 2011. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Gong:2023:RMC

[GWL^{+}23] Jibing Gong, Yao Wan, Ye Liu, Xuewen Li, Yi Zhao, Cheng Wang, Yuting Lin, Xiaohan Fang, Wenzheng Feng, Jingyi Zhang, and Jie Tang. Reinforced MOOCs concept recommendation in heterogeneous information networks. ACM Transactions on the Web (TWEB), 17(3):22:1–22:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580510.

Guo:2016:FEE

[GZC^{+}16] Guangming Guo, Feida Zhu, Enhong Chen, Qi Liu, Le Wu, and Chu Guan. From footprint to evidence: an exploratory study of mining social data for credit scoring. ACM Transactions on the Web (TWEB), 10 (4):22:1–22:??, December 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Guo:2016:NEB

[GZYS16] Guibing Guo, Jie Zhang, and Neil Yorke-Smith. A novel evidence-based Bayesian similarity measure for recommender systems. ACM Transactions on the Web (TWEB), 10(2):8:1– 8:??, May 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

He:2018:EET

[HGC^{+}18] Ming He, Yong Ge, Enhong Chen, Qi Liu, and Xuesong Wang. Exploring the emerging type of comment for online videos: DanMu. ACM Transactions on the Web (TWEB), 12(1):1:1–1:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Hurley:2011:HBP

[HGPS11] John Hurley, Emi Garcia-Palacios, and Sakir Sezer. Host-based P2P flow identification and use in real-time. ACM Transactions on the Web (TWEB), 5(2):7:1–7:??, May 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Hassanpour:2020:IAV

[HHBT20] Masoud Hassanpour, Seyed Amir Hoseinitabatabaei, Payam Barnaghi, and Rahim Tafazolli. Improving the accuracy of the video popularity prediction models through user grouping and video popularity classification. ACM Transactions on the Web (TWEB), 14(1): 4:1–4:28, February 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372499.

Hwang:2016:PPS

[HKH^{+}16] Seung-Won Hwang, Saehoon Kim, Yuxiong He, Sameh Elnikety, and Seungjin Choi. Prediction and predictability for search query acceleration. ACM Transactions on the Web (TWEB), 10(3):19:1–19:??, August 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Hogan:2017:CFI

[Hog17] Aidan Hogan. Canonical forms for isomorphic and equivalent RDF graphs: Algorithms for leaning and labelling blank nodes. ACM Transactions on the Web (TWEB), 11(4):22:1– 22:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Hanson:2013:PWA

[HR13] Vicki L. Hanson and John T. Richards. Progress on Website accessibility? ACM Transactions on the Web (TWEB), 7 (1):2:1–2:??, March 2013. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Huang:2013:UEQ

[Hua13] Xiaodi Huang. UsageQoS: Estimating the QoS of Web services through online user communities. ACM Transactions on the Web (TWEB), 8(1):1:1– 1:??, December 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Haider:2023:CLV

[HY23] Waqar Haider and Yeliz Yesilada. Classification of layout vs. relational tables on the Web: Machine learning with rendered pages. ACM Transactions on the Web (TWEB), 17(1):1:1–1:??, February 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3555349.

Jansen:2007:CES

[Jan07] Bernard J. Jansen. The comparative effectiveness of sponsored and nonsponsored links for Web e-commerce queries. ACM Transactions on the Web (TWEB), 1(1):3:1–3:??, May 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Jackson:2009:PBD

[JBB^{+}09] Collin Jackson, Adam Barth, Andrew Bortz, Weidong Shao, and Dan Boneh. Protecting browsers from DNS rebinding attacks. ACM Transactions on the Web (TWEB), 3(1):2:1– 2:??, January 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Javed:2020:EBD

[JBWR20] Amir Javed, Pete Burnap, Matthew L. Williams, and Omer F. Rana. Emotions behind drive-by download propagation on Twitter. ACM Transactions on the Web (TWEB), 14 (4):16:1–16:26, September 2020. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3408894.

Jiang:2021:RSG

[JCD^{+}21] Wenjun Jiang, Jing Chen, Xiaofei Ding, Jie Wu, Jiawei He, and Guojun Wang. Review summary generation in online systems: Frameworks for supervised and unsupervised scenarios. ACM Transactions on the Web (TWEB), 15(3): 13:1–13:33, July 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448015.

Jurca:2010:RIB

[JGTF10] Radu Jurca, Florent Garcin, Arjun Talwar, and Boi Faltings. Reporting incentives and biases in online review forums. ACM Transactions on the Web (TWEB), 4(2):5:1–5:??, April 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Jeong:2020:DOW

[JJM20] Hyuk-Jin Jeong, Inchang Jeong, and Soo-Mook Moon. Dynamic offloading of Web application execution using snapshot. ACM Transactions on the Web (TWEB), 14(4):15:1– 15:24, September 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3402124.

Jeon:2012:WCP

[JKH^{+}12] Myeongjae Jeon, Youngjae Kim, Jeaho Hwang, Joonwon Lee, and Euiseong Seo. Workload characterization and performance implications of largescale blog servers. ACM Transactions on the Web (TWEB), 6 (4):16:1–16:??, November 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Jain:2023:OLI

[JKS23] Lokesh Jain, Rahul Katarya, and Shelly Sachdeva. Opinion leaders for information diffusion using graph neural network in online social networks. ACM Transactions on the Web (TWEB), 17(2):13:1– 13:??, May 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3580516.

Jones:2024:SWA

[JKWN24] Shawn M. Jones, Martin Klein, Michele C. Weigle, and Michael L. Nelson. Summarizing Web archive corpora via social media storytelling by automatically selecting and visualizing exemplars. ACM Transactions on the Web (TWEB), 18(1):3:1–3:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606030.

Jha:2024:RIA

[JTLM24] Nikhil Jha, Martino Trevisan, Emilio Leonardi, and Marco Mellia. Re-identification attacks against the topics API. ACM Transactions on the Web (TWEB), 18(3):39:1–39:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3675400.

Jha:2022:IPP

[JTVM22] Nikhil Jha, Martino Trevisan, Luca Vassio, and Marco Mellia. The Internet with privacy policies: Measuring the Web upon consent. ACM Transactions on the Web (TWEB), 16(3):15:1–15:??, August 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3555352.

Jia:2018:KGE

[JWJ^{+}18] Yantao Jia, Yuanzhuo Wang, Xiaolong Jin, Hailun Lin, and Xueqi Cheng. Knowledge graph embedding: a locally and temporally adaptive translationbased approach. ACM Transactions on the Web (TWEB), 12(2):8:1–8:??, June 2018. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Jiang:2013:ULI

[JWW^{+}13] Jing Jiang, Christo Wilson, Xiao Wang, Wenpeng Sha, Peng Huang, Yafei Dai, and Ben Y. Zhao. Understanding latent interactions in online social networks. ACM Transactions on the Web (TWEB), 7(4):18:1– 18:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Jiao:2021:SGM

[JXCX21] Simiao Jiao, Zihui Xue, Xiaowei Chen, and Yuedong Xu. Sampling graphlets of multiplex networks: a restricted random walk approach. ACM Transactions on the Web (TWEB), 15 (4):18:1–18:31, July 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3456291.

Koutrika:2008:CST

[KEG^{+}08] Georgia Koutrika, Frans Adjie Effendi, Zolt´ n Gyöngyi, Paul Heymann, and Hector Garcia-Molina. Combating spam in tagging systems: an evaluation. ACM Transactions on the Web (TWEB), 2(4):22:1–22:??, October 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Katzir:2015:ECC

[KH15] Liran Katzir and Stephen J.

Hardiman. Estimating clustering coefficients and size of social networks via random walk. ACM Transactions on the Web (TWEB), 9(4):19:1–19:??, October 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kilic:2023:PSO

[KI23] Yasir Kilic and Ali Inan. Privacy scoring over OSNs: Shared data granularity as a latent dimension. ACM Transactions on the Web (TWEB), 17(4):30:1– 30:??, November 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604909.

Kanza:2017:LBD

[KKSS17] Yaron Kanza, Elad Kravi, Eliyahu Safra, and Yehoshua Sagiv. Location-based distance measures for geosocial similarity. ACM Transactions on the Web (TWEB), 11(3):17:1– 17:??, July 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kiciman:2010:APR

[KL10] Emre Kiciman and Benjamin Livshits. AjaxScope: a platform for remotely monitoring the client-side behavior of Web 2.0 applications. ACM Transactions on the Web (TWEB), 4 (4):13:1–13:??, September 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kang:2017:EMA

[KL17a] Jeon-Hyung Kang and Kristina Lerman. Effort mediates access to information in online social networks. ACM Transactions on the Web (TWEB), 11(1):3:1– 3:??, March 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Koutrika:2017:SWP

[KL17b] Georgia Koutrika and Qian Lin. A study of Web print: What people print in the digital era. ACM Transactions on the Web (TWEB), 11(4):23:1– 23:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kumar:2024:CEL

[KMSB24] Mukesh Kumar, Shivansh Mishra, Shashank Sheshar Singh, and Bhaskar Biswas. Community-enhanced link prediction in dynamic networks. ACM Transactions on the Web (TWEB), 18(2):24:1– 24:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3580513.

Kwasnikowska:2015:FAO

[KMV15] Natalia Kwasnikowska, Luc Moreau, and Jan Van Den Bussche. A formal account of the open provenance model. ACM Transactions on the Web (TWEB), 9(2):10:1–10:??, May 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kyusakov:2014:EFE

[KPED14] Rumen Kyusakov, Pablo Puñal Pereira, Jens Eliasson, and Jerker Delsing. EXIP: a framework for embedded Web development. ACM Transactions on the Web (TWEB), 8(4):23:1– 23:??, October 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Karanatsiou:2022:MTB

[KSG^{+}22] Dimitra Karanatsiou, Pavlos Sermpezis, Dritjon Gruda, Konstantinos Kafetsios, Ilias Dimitriadis, and Athena Vakali. My tweets bring all the traits to the yard: Predicting personality and relational traits in online social networks. ACM Transactions on the Web (TWEB), 16 (2):10:1–10:26, May 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3523749.

Koutsonikola:2011:CDL

[KV11] Vassiliki Koutsonikola and Athena Vakali. A clusteringdriven LDAP framework. ACM Transactions on the Web (TWEB), 5(3):12:1–12:??, July 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Kitzler:2023:DDF

[KVSH23] Stefan Kitzler, Friedhelm Victor, Pietro Saggese, and Bernhard Haslhofer. Disentangling decentralized finance (DeFi) compositions. ACM Transactions on the Web (TWEB), 17(2):10:1–10:??, May 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532857.

Kaldeli:2013:CWS

[KWLA13] Eirini Kaldeli, Ehsan Ullah Warriach, Alexander Lazovik, and Marco Aiello. Coordinating the web of services for a smart home. ACM Transactions on the Web (TWEB), 7(2):10:1– 10:??, May 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Leskovec:2007:DVM

[LAH07] Jure Leskovec, Lada A. Adamic, and Bernardo A. Huberman. The dynamics of viral marketing. ACM Transactions on the Web (TWEB), 1(1):5:1– 5:??, May 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Laperdrix:2020:BFS

[LBBA20] Pierre Laperdrix, Nataliia Bielova, Benoit Baudry, and Gildas Avoine. Browser fingerprinting: a survey. ACM Transactions on the Web (TWEB), 14(2):8:1–8:33, April 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386040.

Liu:2012:IWS

[LCK^{+}12] Yiqun Liu, Fei Chen, Weize Kong, Huijia Yu, Min Zhang, Shaoping Ma, and Liyun Ru. Identifying Web spam with the wisdom of the crowds. ACM Transactions on the Web (TWEB), 6(1):2:1–2:??, March 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Liu:2019:FPS

[LCLQ19] Daxin Liu, Gong Cheng, Qingxia Liu, and Yuzhong Qu. Fast and practical snippet generation for RDF datasets. ACM Transactions on the Web (TWEB), 13(4): 19:1–19:??, December 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3365575.

Li:2020:TAS

[LCZ^{+}20] Guohui Li, Qi Chen, Bolong Zheng, Nguyen Quoc Viet Hung, Pan Zhou, and Guanfeng Liu. Time-aspect-sentiment recommendation models based on novel similarity measure methods. ACM Transactions on the Web (TWEB), 14(2): 5:1–5:26, April 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3375548.

Lee:2015:DPM

[Lee15] Sihyung Lee. Detection of political manipulation in online communities through measures of effort and collaboration. ACM Transactions on the Web (TWEB), 9(3):16:1–16:??, June 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Li:2024:HCP

[LFAH24] Lingyao Li, Lizhou Fan, Shubham Atreja, and Libby Hemphill. “HOT” ChatGPT: The promise of ChatGPT in detecting and discriminating hateful, offensive, and toxic comments on social media. ACM Transactions on the Web (TWEB), 18(2): 30:1–30:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3643829.

Lee:2013:SCA

[LHJL13] Jung-Hyun Lee, Jongwoo Ha, Jin-Yong Jung, and Sangkeun Lee. Semantic contextual advertising based on the open directory project. ACM Transactions on the Web (TWEB), 7(4):24:1– 24:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Liao:2013:VAC

[LJP^{+}13] Zhen Liao, Daxin Jiang, Jian Pei, Yalou Huang, Enhong Chen, Huanhuan Cao, and Hang Li. A vlHMM approach to context-aware search. ACM Transactions on the Web (TWEB), 7(4):22:1–22:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Lin:2024:ARI

[LLA^{+}24] Fuqi Lin, Xuan Lu, Wei Ai, Huoran Li, Yun Ma, Yulian Yang, Hongfei Deng, Qingxiang Wang, Qiaozhu Mei, and Xuanzhe Liu. Adoption of recurrent innovations: a large-scale case study on mobile app updates. ACM Transactions on the Web (TWEB), 18(1):13:1– 13:??, February 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3626189.

Liu:2013:SCB

[LLM13] Liwei Liu, Freddy Lecue, and Nikolay Mehandjiev. Semantic content-based recommendation of software services using context. ACM Transactions on the Web (TWEB), 7(3):17:1–17:??, September 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Liu:2018:RCW

[LLSL18] Yining Liu, Yong Liu, Yanming Shen, and Keqiu Li. Recommendation in a changing world: Exploiting temporal dynamics in ratings and reviews. ACM Transactions on the Web (TWEB), 12(1):3:1–3:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Lauw:2012:QLO

[LLW12] Hady W. Lauw, Ee-Peng Lim, and Ke Wang. Quality and leniency in online collaborative rating systems. ACM Transactions on the Web (TWEB), 6 (1):4:1–4:??, March 2012. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Li:2023:TIU

[LLW^{+}23] Qian Li, Jianxin Li, Lihong Wang, Cheng Ji, Yiming Hei, Jiawei Sheng, Qingyun Sun, Shan Xue, and Pengtao Xie. Type information utilized event detection via multi-channel GNNs in electrical power systems. ACM Transactions on the Web (TWEB), 17(3):20:1– 20:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3577031.

Lee:2009:ISB

[LLWL09] Hsin-Tsang Lee, Derek Leonard, Xiaoming Wang, and Dmitri Loguinov. IRLbot: Scaling to 6 billion pages and beyond. ACM Transactions on the Web (TWEB), 3(3):8:1–8:??, June 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Li:2010:DSO

[LMJ10] Guoli Li, Vinod Muthusamy, and Hans-Arno Jacobsen. A distributed service-oriented architecture for business process execution. ACM Transactions on the Web (TWEB), 4(1):2:1– 2:??, January 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Li:2024:MLS

[LMLC24] Qingqing Li, Huifang Ma, Zhixin Li, and Liang Chang. Multiresolution local spectral attributed community search. ACM Transactions on the Web (TWEB), 18(1):9:1–9:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624580.

Lin:2008:DST

[LSC^{+}08] Yu-Ru Lin, Hari Sundaram, Yun Chi, Junichi Tatemura, and Belle L. Tseng. Detecting splogs via temporal dynamics using self-similarity analysis. ACM Transactions on the Web (TWEB), 2(1):4:1–4:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Liu:2024:BFG

[LTS^{+}24] Mingyi Liu, Zhiying Tu, Tonghua Su, Xianzhi Wang, Xiaofei Xu, and Zhongjie Wang. BehaviorNet: a finegrained behavior-aware network for dynamic link prediction. ACM Transactions on the Web (TWEB), 18(2):25:1– 25:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3580514.

Leiva:2013:WBB

[LV13] Luis A. Leiva and Roberto Vivó. Web browsing behavior analysis and interactive hypervideo. ACM Transactions on the Web (TWEB), 7(4):20:1–20:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Liu:2023:CCU

[LYZ23] Zhenfang Liu, Jianxiong Ye, and Zhaonian Zou. Closeness centrality on uncertain graphs. ACM Transactions on the Web (TWEB), 17(4):29:1– 29:??, November 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3604912.

Liu:2022:FHU

[LZKN22] Bang Liu, Hanlin Zhang, Linglong Kong, and Di Niu. Factorizing historical user actions for next-day purchase prediction. ACM Transactions on the Web (TWEB), 16(1):1:1– 1:26, February 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3468227.

Luo:2024:SIM

[LZX^{+}24] Pengfei Luo, Xi Zhu, Tong Xu, Yi Zheng, and Enhong Chen. Semantic interaction matching network for few-shot knowledge graph completion. ACM Transactions on the Web (TWEB), 18 (2):20:1–20:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3589557.

Margaritis:2014:ITI

[MA14] Giorgos Margaritis and Stergios V. Anastasiadis. Incremental text indexing for fast diskbased search. ACM Transactions on the Web (TWEB), 8 (3):16:1–16:??, June 2014. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Magno:2022:MIO

[MA22] Gabriel Magno and Virgilio Almeida. Measuring international online human values with word embeddings. ACM Transactions on the Web (TWEB), 16(2):9:1–9:38, May 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3501306.

Mitra:2011:CWB

[MAY^{+}11] Siddharth Mitra, Mayank Agrawal, Amit Yadav, Niklas Carlsson, Derek Eager, and Anirban Mahanti. Characterizing Web-based video sharing workloads. ACM Transactions on the Web (TWEB), 5(2):8:1– 8:??, May 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Manta-Caro:2018:MSW

[MCFL18] Cristyan Manta-Caro and Juan M. Fernández-Luna. Modeling and simulating the Web of Things from an information retrieval perspective. ACM Transactions on the Web (TWEB), 12(1):6:1– 6:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Manica:2019:CUH

[MDG19] Edimar Manica, Carina Friedrich Dorneles, and Renata Galante. Combining URL and HTML features for entity discovery in the Web. ACM Transactions on the Web (TWEB), 13 (4):20:1–20:??, December 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3365574.

Mistry:2021:SLB

[MFB21] Sajib Mistry, Sheik Mohammad Mostakim Fattah, and Athman Bouguettaya. Sequential learning-based IaaS composition. ACM Transactions on the Web (TWEB), 15(3): 15:1–15:37, July 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3452332.

Ma:2020:RTC

[MHG^{+}20] Yun Ma, Ziniu Hu, Diandian Gu, Li Zhou, Qiaozhu Mei, Gang Huang, and Xuanzhe Liu. Roaming through the castle tunnels: an empirical analysis of inter-app navigation of Android apps. ACM Transactions on the Web (TWEB), 14 (3):14:1–14:24, July 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3395050.

Mukherjee:2017:ISV

[MJ17] Partha Mukherjee and Bernard J.

Jansen. Information sharing by viewers via second screens for in-real-life events. ACM Transactions on the Web (TWEB), 11 (1):1:1–1:??, March 2017. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Miliaraki:2012:FDS

[MK12] Iris Miliaraki and Manolis Koubarakis. FoXtrot: Distributed structural and value XML filtering. ACM Transactions on the Web (TWEB), 6 (3):12:1–12:??, September 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Moon:2022:MME

[MKH22] Taegeun Moon, Hyoungshick Kim, and Sangwon Hyun. Mutexion: Mutually exclusive compression system for mitigating compression side-channel attacks. ACM Transactions on the Web (TWEB), 16(4):21:1– 21:??, November 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3532850.

Mohan:2007:SPC

[MKR07] Bharath Kumar Mohan, Benjamin J. Keller, and Naren Ramakrishnan. Scouts, promoters, and connectors: The roles of ratings in nearestneighbor collaborative filtering. ACM Transactions on the Web (TWEB), 1(2):8:1–8:??, August 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Mula:2018:FBE

[ML18] Wojciech Muła and Daniel Lemire. Faster Base64 encoding and decoding using AVX2 instructions. ACM Transactions on the Web (TWEB), 12(3): 20:1–20:??, July 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Merhav:2012:EIN

[MMB^{+}12] Yuval Merhav, Filipe Mesquita, Denilson Barbosa, Wai Gen Yee, and Ophir Frieder. Extracting information networks from the blogosphere. ACM Transactions on the Web (TWEB), 6(3):11:1–11:??, September 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Marriott:2013:HAT

[MMH13] Kim Marriott, Peter Moulder, and Nathan Hurst. HTML automatic table layout. ACM Transactions on the Web (TWEB), 7 (1):4:1–4:??, March 2013. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Marszalkowski:2016:ASC

[MMMD16] Jakub Marszalkowski, Jan Mizgajski, Dariusz Mokwa, and Maciej Drozdowski. Analysis and solution of CSS-sprite packing problem. ACM Transactions on the Web (TWEB), 10(1):1:1– 1:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Maceiras:2024:KTC

[MNB^{+}24] Maël Maceiras, Kavous Salehzadeh Niksirat, Gaël Bernard, Benoit Garbinato, Mauro Cherubini, Mathias Humbert, and Kévin Huguenin. Know their customers: an empirical study of online account enumeration attacks. ACM Transactions on the Web (TWEB), 18(3):37:1– 37:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3664201.

Milton:2023:UES

[MP23] Ashlee Milton and Maria Soledad Pera. Into the unknown: Exploration of search engines’ responses to users with depression and anxiety. ACM Transactions on the Web (TWEB), 17 (4):25:1–25:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3580283.

Mazumdar:2020:CSP

[MPB20] Pramit Mazumdar, Bidyut Kr. Patra, and Korra Sathya Babu. Cold-start point-of-interest recommendation through crowdsourcing. ACM Transactions on the Web (TWEB), 14(4):19:1– 19:36, September 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3407182.

Montali:2010:DSV

[MPvdA^{+}10] Marco Montali, Maja Pesic, Wil M. P. van der Aalst, Federico Chesani, Paola Mello, and Sergio Storari. Declarative specification and verification of service choreographies. ACM Transactions on the Web (TWEB), 4 (1):3:1–3:??, January 2010. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Malak:2010:MWQ

[MSBB10] Ghazwa Malak, Houari Sahraoui, Linda Badri, and Mourad Badri. Modeling Web quality using a probabilistic approach: an empirical validation. ACM Transactions on the Web (TWEB), 4(3):9:1–9:??, July 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Matsubara:2017:NDI

[MSP^{+}17] Yasuko Matsubara, Yasushi Sakurai, B. Aditya Prakash, Lei Li, and Christos Faloutsos. Nonlinear dynamics of information diffusion in social networks. ACM Transactions on the Web (TWEB), 11(2):11:1– 11:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Minervini:2018:AKP

[MTDF18] Pasquale Minervini, Volker Tresp, Claudia D’amato, and Nicola Fanizzi. Adaptive knowledge propagation in Web ontologies. ACM Transactions on the Web (TWEB), 12(1):2:1– 2:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Mesbah:2012:CAB

[MvDL12] Ali Mesbah, Arie van Deursen, and Stefan Lenselink. Crawling Ajax-based Web applications through dynamic analysis of user interface state changes. ACM Transactions on the Web (TWEB), 6(1):3:1–3:??, March 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ma:2024:CCA

[MWHW24] Yutao Ma, Zesheng Wang, Liwei Huang, and Jian Wang. CLHHN: Category-aware lossless heterogeneous hypergraph neural network for session-based recommendation. ACM Transactions on the Web (TWEB), 18 (1):12:1–12:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3626569.

Mittos:2020:AGT

[MZBD20] Alexandros Mittos, Savvas Zannettou, Jeremy Blackburn, and Emiliano De Cristofaro. Analyzing genetic testing discourse on the Web through the lens of Twitter, Reddit, and 4chan. ACM Transactions on the Web (TWEB), 14 (4):17:1–17:38, September 2020. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3404994.

Naini:2016:SEW

[NAS16] Kaweh Djafari Naini, Ismail Sengor Altingovde, and Wolf Siberski. Scalable and efficient Web search result diversification. ACM Transactions on the Web (TWEB), 10(3):15:1– 15:??, August 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Najork:2008:ISS

[ND08] Marc Najork and Brian D. Davison. Introduction to special section on adversarial issues in Web search. ACM Transactions on the Web (TWEB), 2(1):1:1– 1:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Nguyen:2013:FWT

[NKTP13] Cam-Tu Nguyen, Natsuda Kaothanthong, Takeshi Tokuyama, and Xuan-Hieu Phan. A feature-word-topic model for image annotation and retrieval. ACM Transactions on the Web (TWEB), 7(3):12:1–12:??, September 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Nelson:2022:QCT

[NRS^{+}22] Michael Nelson, Sridhar Radhakrishnan, Chandra Sekharan, Amlan Chatterjee, and Sudhindra Gopal Krishna. Queryable compression on time-evolving Web and social networks with streaming. ACM Transactions on the Web (TWEB), 16(2): 6:1–6:21, May 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3495012.

Ozcan:2013:SCH

[OACU13] Rifat Ozcan, Ismail Sengor Altingovde, B. Barla Cambazoglu, and Özgür Ulusoy. Second chance: a hybrid approach for dynamic result caching and prefetching in search engines. ACM Transactions on the Web (TWEB), 8(1):3:1–3:??, December 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Ozcan:2011:CAS

[OAU11] Rifat Ozcan, Ismail Sengor Altingovde, and Özgür Ulusoy. Cost-aware strategies for query result caching in Web search engines. ACM Transactions on the Web (TWEB), 5(2):9:1– 9:??, May 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Oswald:2022:SIA

[OSB22] C. Oswald, Sona Elza Simon, and Arnab Bhattacharya. SpotSpam: Intention analysisdriven SMS spam detection using BERT embeddings. ACM Transactions on the Web (TWEB), 16(3):14:1–14:??, August 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3538491.

Parra-Arnau:2017:MBT

[PAAC17] Javier Parra-Arnau, Jagdish Prasad Achara, and Claude Castelluccia. MyAdChoices: Bringing transparency and control to online advertising. ACM Transactions on the Web (TWEB), 11 (1):7:1–7:??, March 2017. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Parikh:2021:CSM

[PAC^{+}21] Pulkit Parikh, Harika Abburi, Niyati Chhaya, Manish Gupta, and Vasudeva Varma. Categorizing sexism and misogyny through neural approaches. ACM Transactions on the Web (TWEB), 15(4): 17:1–17:31, July 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3457189.

Park:2018:LLB

[PAL18] Sangkeun Park, Mark S. Ackerman, and Uichin Lee. Localness of location-based knowledge sharing: a study of Naver KiN “Here”. ACM Transactions on the Web (TWEB), 12 (3):16:1–16:??, July 2018. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Pugliese:2014:EMM

[PBSO14] Andrea Pugliese, Matthias Bröcheler, V. S. Subrahmanian, and Michael Ovelgönne. Efficient MultiView maintenance under insertion in huge social networks. ACM Transactions on the Web (TWEB), 8(2):10:1– 10:??, March 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Pajola:2023:NRH

[PCCS23] Luca Pajola, Dongkai Chen, Mauro Conti, and V. S. Subrahmanian. A novel review helpfulness measure based on the user-review-item paradigm. ACM Transactions on the Web (TWEB), 17(4):27:1– 27:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3585280.

Piccardi:2023:LSC

[PGAW23] Tiziano Piccardi, Martin Gerlach, Akhil Arora, and Robert West. A large-scale characterization of how readers browse Wikipedia. ACM Transactions on the Web (TWEB), 17(2): 11:1–11:??, May 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3580318.

Panagopoulos:2017:MER

[PKT17] A. Panagopoulos, E. Koutrouli, and A. Tsalgatidou. Modeling and evaluating a robust feedback-based reputation system for e-commerce platforms. ACM Transactions on the Web (TWEB), 11(3):18:1– 18:??, July 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Park:2018:WSD

[PMGO18] Souneil Park, Aleksandar Matic, Kamini Garg, and Nuria Oliver. When simpler data does not imply less information: a study of user profiling scenarios with constrained view of mobile HTTP(S) traffic. ACM Transactions on the Web (TWEB), 12(2):9:1–9:??, June 2018. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Paci:2011:AAC

[PMOB11] Federica Paci, Massimo Mecella, Mourad Ouzzani, and Elisa Bertino. ACConv – an access control model for conversational Web services. ACM Transactions on the Web (TWEB), 5(3):13:1–13:??, July 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Poddar:2024:MQC

[PMS^{+}24] Soham Poddar, Rajdeep Mukherjee, Azlaan Samad, Niloy Ganguly, and Saptarshi Ghosh. MuLX-QA: Classifying multilabels and extracting rationale spans in social media posts. ACM Transactions on the Web (TWEB), 18(3):35:1– 35:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3653303.

Parisi:2018:TKU

[PPPS18] Francesco Parisi, Noseong Park, Andrea Pugliese, and V. S. Subrahmanian. Topk user-defined vertex scoring queries in edge-labeled graph databases. ACM Transactions on the Web (TWEB), 12 (4):21:1–21:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Poblete:2010:PPQ

[PSBY10] Barbara Poblete, Myra Spiliopoulou, and Ricardo Baeza-Yates. Privacypreserving query log mining for business confidentiality protection. ACM Transactions on the Web (TWEB), 4(3):10:1–10:??, July 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Pathak:2023:UCR

[PSP23] Royal Pathak, Francesca Spezzano, and Maria Soledad Pera. Understanding the contribution of recommendation algorithms on misinformation recommendation and misinformation dissemination on social networks. ACM Transactions on the Web (TWEB), 17(4):35:1– 35:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3616088.

Pilioura:2009:UPD

[PT09] Thomi Pilioura and Aphrodite Tsalgatidou. Unified publication and discovery of semantic Web services. ACM Transactions on the Web (TWEB), 3 (3):11:1–11:??, June 2009. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Poiitis:2022:ADM

[PVK22] Marinos Poiitis, Athena Vakali, and Nicolas Kourtellis. On the aggression diffusion modeling and minimization in Twitter. ACM Transactions on the Web (TWEB), 16(1):5:1– 5:24, February 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3486218.

Perino:2019:LTM

[PVS19] Diego Perino, Matteo Varvello, and Claudio Soriente. Longterm measurement and analysis of the free proxy ecosystem. ACM Transactions on the Web (TWEB), 13(4):18:1– 18:??, December 2019. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3360695.

Paul:2016:SBC

[PWH16] Michael J. Paul, Ryen W.

White, and Eric Horvitz. Search and breast cancer: On episodic shifts of attention over life histories of an illness. ACM Transactions on the Web (TWEB), 10 (2):13:1–13:??, May 2016. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Peng:2023:ISI

[PYWY23] Hao Peng, Jian Yang, Jia Wu, and Philip S. Yu. Introduction to the special issue on advanced graph mining on the Web: Theory, algorithms, and applications: Part 1. ACM Transactions on the Web (TWEB), 17(3):14:1–14:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579360.

Peng:2024:ISI

[PYWY24] Hao Peng, Jian Yang, Jia Wu, and Philip S. Yu. Introduction to the special issue on advanced graph mining on the Web: Theory, algorithms, and applications: Part 2. ACM Transactions on the Web (TWEB), 18 (2):16:1–16:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3631941.

Qian:2014:FTD

[QA14] Yi Qian and Sibel Adali. Foundations of trust and distrust in networks: Extended structural balance theory. ACM Transactions on the Web (TWEB), 8 (3):13:1–13:??, June 2014. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Quarteroni:2013:BKA

[QBC13] Silvia Quarteroni, Marco Brambilla, and Stefano Ceri. A bottom-up, knowledge-aware approach to integrating and querying Web data services. ACM Transactions on the Web (TWEB), 7(4):19:1–19:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Qian:2022:PVR

[QRD^{+}22] Xin Qian, Ryan A. Rossi, Fan Du, Sungchul Kim, Eunyee Koh, Sana Malik, Tak Yeon Lee, and Nesreen K. Ahmed. Personalized visualization recommendation. ACM Transactions on the Web (TWEB), 16(3):11:1–11:??, August 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3538703.

Qiao:2024:DCS

[QWW^{+}24] Ziyue Qiao, Pengyang Wang, Pengfei Wang, Zhiyuan Ning, Yanjie Fu, Yi Du, Yuanchun Zhou, Jianqiang Huang, Xian-Sheng Hua, and Hui Xiong. A dual-channel semi-supervised learning framework on graphs via knowledge transfer and meta-learning. ACM Transactions on the Web (TWEB), 18(2):18:1–18:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3577033.

Ryu:2008:SDE

[RCS^{+}08] Seung Hwan Ryu, Fabio Casati, Halvard Skogsrud, Boualem Benatallah, and Régis Saint-Paul. Supporting the dynamic evolution of Web service protocols in service-oriented architectures. ACM Transactions on the Web (TWEB), 2(2):13:1– 13:??, April 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rogers:2007:EPB

[RDJS07] Alex Rogers, Esther David, Nicholas R. Jennings, and Jeremy Schiff. The effects of proxy bidding and minimum bid increments within eBay auctions. ACM Transactions on the Web (TWEB), 1(2):9:1–9:??, August 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Reay:2009:LSE

[RDM09] Ian Reay, Scott Dick, and James Miller. A large-scale empirical study of P3P privacy policies: Stated actions vs. legal obligations. ACM Transactions on the Web (TWEB), 3(2):6:1– 6:??, April 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rieger:2024:NMC

[RDTT24] Alisa Rieger, Tim Draws, Mariët Theune, and Nava Tintarev. Nudges to mitigate confirmation bias during Web search on debated topics: Support vs. manipulation. ACM Transactions on the Web (TWEB), 18(2):27:1– 27:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3635034.

Reis:2007:BVD

[RDW^{+}07] Charles Reis, John Dunagan, Helen J. Wang, Opher Dubrovsky, and Saher Esmeir. BrowserShield: Vulnerabilitydriven filtering of dynamic HTML. ACM Transactions on the Web (TWEB), 1(3):11:1– 11:??, September 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rafalak:2016:WCC

[RDW^{+}16] Maria Rafalak, Dominik Deja, Adam Wierzbicki, Radoslaw Nielek, and Michal Kakol. Web content classification using distributions of subjective quality evaluations. ACM Transactions on the Web (TWEB), 10 (4):21:1–21:??, December 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rojas-Galeano:2017:OOO

[RG17]  Sergio Rojas-Galeano.  On

obstructing obscenity obfuscation. ACM Transactions on the Web (TWEB), 11(2):12:1– 12:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rudra:2018:ESS

[RGGG18] Koustav Rudra, Niloy Ganguly, Pawan Goyal, and Saptarshi Ghosh. Extracting and summarizing situational information from the Twitter social media during disasters. ACM Transactions on the Web (TWEB), 12 (3):17:1–17:??, July 2018. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Redmiles:2019:NPW

[RH19] Elissa M. Redmiles and Eszter Hargittai. New phone, who dis? Modeling millennials’ backup behavior. ACM Transactions on the Web (TWEB), 13(1):4:1–4:??, February 2019.

CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rahman:2017:AAC

[RHLC17] M. Rezaur Rahman, Jinyoung Han, Yong Jae Lee, and Chen-Nee Chuah. Analyzing the adoption and cascading process of OSN-based gifting applications: an empirical study. ACM Transactions on the Web (TWEB), 11(2):10:1– 10:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Robu:2009:ECS

[RHS09] Valentin Robu, Harry Halpin, and Hana Shepherd. Emergence of consensus and shared vocabularies in collaborative tagging systems. ACM Transactions on the Web (TWEB), 3(4):14:1– 14:??, September 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Richardson:2008:LAW

[Ric08] Matthew Richardson. Learning about the world through longterm query logs. ACM Transactions on the Web (TWEB), 2(4):21:1–21:??, October 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Raponi:2022:FNP

[RKOD22] Simone Raponi, Zeinab Khalifa, Gabriele Oligeri, and Roberto Di Pietro. Fake news propagation: a review of epidemic models, datasets, and insights. ACM Transactions on the Web (TWEB), 16(3): 12:1–12:??, August 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522756.

Rattenbury:2009:MEP

[RN09] Tye Rattenbury and Mor Naaman. Methods for extracting place semantics from Flickr tags. ACM Transactions on the Web (TWEB), 3(1):1:1– 1:??, January 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Rocha:2017:LPL

[RP17] André Rocha and Cássio Prazeres. LDoW–PaN: Linked data on the Web-presentation and navigation. ACM Transactions on the Web (TWEB), 11 (4):20:1–20:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Riedl:2011:ISI

[RS11] John Riedl and Barry Smyth.

Introduction to special issue on recommender systems. ACM Transactions on the Web (TWEB), 5(1):1:1–1:??, February 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Roy:2021:IAI

[RSC^{+}21] Soumyadeep Roy, Shamik Sural, Niyati Chhaya, Anandhavelu Natarajan, and Niloy Ganguly. An integrated approach for improving brand consistency of Web content: Modeling, analysis, and recommendation. ACM Transactions on the Web (TWEB), 15(2):9:1–9:25, May 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450445.

Romero:2019:SNU

[RUK19] Daniel M. Romero, Brian Uzzi, and Jon Kleinberg. Social networks under stress: Specialized team roles and their communication structure. ACM Transactions on the Web (TWEB), 13(1):6:1–6:??, February 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sharma:2024:URP

[SAS24] Trishie Sharma, Rachit Agarwal, and Sandeep Kumar Shukla. Understanding rug pulls: an in-depth behavioral analysis of fraudulent NFT creators. ACM Transactions on the Web (TWEB), 18(1):8:1– 8:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3623376.

Srba:2016:CSC

[SB16] Ivan Srba and Maria Bielikova. A comprehensive survey and classification of approaches for community question answering. ACM Transactions on the Web (TWEB), 10(3):18:1–18:??, August 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Shah:2024:EIA

[SB24] Chirag Shah and Emily M. Bender. Envisioning information access systems: What makes for good tools and a healthy Web? ACM Transactions on the Web (TWEB), 18(3): 33:1–33:??, August 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649468.

Squicciarini:2017:TAO

[SCB17] Anna Squicciarini, Cornelia Caragea, and Rahul Balakavi. Toward automated online photo privacy. ACM Transactions on the Web (TWEB), 11(1):2:1– 2:??, March 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Siersdorfer:2014:AMC

[SCS^{+}14] Stefan Siersdorfer, Sergiu Chelaru, Jose San Pedro, Ismail Sengor Altingovde, and Wolfgang Nejdl. Analyzing and mining comments and comment ratings on the social Web. ACM Transactions on the Web (TWEB), 8 (3):17:1–17:??, June 2014. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Swaminathan:2010:RRM

[SCW^{+}10] Ashwin Swaminathan, Renan G. Cattelan, Ydo Wexler, Cherian V. Mathew, and Darko Kirovski. Relating reputation and money in online markets. ACM Transactions on the Web (TWEB), 4(4):17:1–17:??, September 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Soi:2014:CDC

[SDC14] Stefano Soi, Florian Daniel, and Fabio Casati. Conceptual development of custom, domain-specific mashup platforms. ACM Transactions on the Web (TWEB), 8(3):14:1– 14:??, June 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sun:2023:GDR

[SDG^{+}23] Li Sun, Yang Du, Shuai Gao, Junda Ye, Feiyang Wang, Fuxin Ren, Mingchen Liang, Yue Wang, and Shuhai Wang. GroupAligner: a deep reinforcement learning with domain adaptation for social group alignment. ACM Transactions on the Web (TWEB), 17(3):17:1–17:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580509.

Schafer:2008:EFA

[SDN08] Michael Schäfer, Peter Dolog, and Wolfgang Nejdl. An environment for flexible advanced compensations of Web service transactions. ACM Transactions on the Web (TWEB), 2 (2):14:1–14:??, April 2008. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Shrestha:2024:JCE

[SDSS24] Anu Shrestha, Jason Duran, Francesca Spezzano, and Edoardo Serra. Joint credibility estimation of news, user, and publisher via role-relational graph convolutional networks. ACM Transactions on the Web (TWEB), 18(1):6:1–6:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617418.

Sun:2023:ICW

[SFJ^{+}23] Chang-Ai Sun, An Fu, Jingting Jia, Meng Li, and Jun Han. Improving conformance of Web services: a constraint-based model-driven approach. ACM Transactions on the Web (TWEB), 17(2): 12:1–12:??, May 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3580515.

Shah:2020:OMA

[SGJC20] Ankit Shah, Rajesh Ganesan, Sushil Jajodia, and Hasan Cam. An outsourcing model for alert analysis in a cybersecurity operations center. ACM Transactions on the Web (TWEB), 14(1):2:1–2:22, February 2020. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372498.

Singer:2017:BMC

[SHHS17] Philipp Singer, Denis Helic, Andreas Hotho, and Markus Strohmaier. A Bayesian method for comparing hypotheses about human trails. ACM Transactions on the Web (TWEB), 11 (3):14:1–14:??, July 2017. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Su:2014:HIY

[SHKK14] Ao-Jan Su, Y. Charlie Hu, Aleksandar Kuzmanovic, and Cheng-Kok Koh. How to improve your search engine ranking: Myths and reality. ACM Transactions on the Web (TWEB), 8(2):8:1–8:??, March 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Skenderi:2024:IFD

[SHLS24] Erjon Skenderi, Jukka Huhtamäki, Salla-Maaria Laaksonen, and Kostas Stefanidis. IN-CEPT: a framework for duplicate posts classification with combined text representations. ACM Transactions on the Web (TWEB), 18(3):40:1–40:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3677322.

Srivatsa:2008:MAL

[SIYL08] Mudhakar Srivatsa, Arun Iyengar, Jian Yin, and Ling Liu. Mitigating application-level denial of service attacks on Web servers: a client-transparent approach. ACM Transactions on the Web (TWEB), 2(3):15:1– 15:??, July 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sirivianos:2014:LSF

[SKGY14] Michael Sirivianos, Kyungbaek Kim, Jian Wei Gan, and Xiaowei Yang. Leveraging social feedback to verify online identity claims. ACM Transactions on the Web (TWEB), 8(2):9:1– 9:??, March 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Su:2015:RRT

[SLL^{+}15] Zhiyuan Su, Ling Liu, Mingchu Li, Xinxin Fan, and Yang Zhou. Reliable and resilient trust management in distributed service provision networks. ACM Transactions on the Web (TWEB), 9(3):14:1– 14:??, June 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sherkat:2013:ETS

[SLM13] Reza Sherkat, Jing Li, and Nikos Mamoulis. Efficient time-stamped event sequence anonymization. ACM Transactions on the Web (TWEB), 8(1):4:1–4:??, December 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sigg:2019:EUP

[SLP^{+}19] Stephan Sigg, Eemil Lagerspetz, Ella Peltonen, Petteri Nurmi, and Sasu Tarkoma. Exploiting usage to predict instantaneous app popularity: Trend filters and retention rates. ACM Transactions on the Web (TWEB), 13(2):13:1–13:??, April 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3199677.

Serrano:2007:DSW

[SMB^{+}07] M. Ángeles Serrano, Ana Maguitman, Marián Boguñá, Santo Fortunato, and Alessandro Vespignani. Decoding the structure of the WWW: a comparative analysis of Web crawls. ACM Transactions on the Web (TWEB), 1(2):10:1–10:??, August 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sun:2007:MDW

[SMRM07] Zan Sun, Jalal Mahmud, I. V. Ramakrishnan, and Saikat Mukherjee. Modeldirected Web transactions under constrained modalities. ACM Transactions on the Web (TWEB), 1(3):12:1–12:??, September 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sheshbolouki:2023:SES

[SÖ23] Aida Sheshbolouki and M. Tamer Özsu. sGrow: Explaining the scale-invariant strength assortativity of streaming butterflies. ACM Transactions on the Web (TWEB), 17(3):24:1– 24:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3572408.

Sun:2023:PTA

[SQL^{+}23] Ke Sun, Tieyun Qian, Chenliang Li, Xuan Ma, Qing Li, Ming Zhong, Yuanyuan Zhu, and Mengchi Liu. Pretraining across different cities for next POI recommendation. ACM Transactions on the Web (TWEB), 17(4):31:1– 31:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3605554.

Squicciarini:2022:EUG

[SRG^{+}22] Anna Squicciarini, Sarah Rajtmajer, Yang Gao, Justin Semonsen, Andrew Belmonte, and Pratik Agarwal. An extended ultimatum game for multi-party access control in social networks. ACM Transactions on the Web (TWEB), 16(3):13:1– 13:??, August 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3555351.

Sharman:2007:CAD

[SRRG07] Raj Sharman, Shiva Shankar Ramanna, Ram Ramesh, and Ram Gopal. Cache architecture for on-demand streaming on the Web. ACM Transactions on the Web (TWEB), 1(3):13:1–13:??, September 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Seneviratne:2017:SMA

[SSK^{+}17] Suranga Seneviratne, Aruna Seneviratne, Mohamed Ali Kaafar, Anirban Mahanti, and Prasant Mohapatra. Spam mobile apps: Characteristics, detection, and in the wild analysis. ACM Transactions on the Web (TWEB), 11(1):4:1–4:??, March 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Singh:2009:SSO

[SSL09] Aameek Singh, Mudhakar Srivatsa, and Ling Liu. Searchas-a-service: Outsourced search over outsourced storage. ACM Transactions on the Web (TWEB), 3(4):13:1–13:??, September 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sariyuce:2017:NDI

[SSPÇ17] Ahmet Erdem Sariyüce, C. Seshadhri, Ali Pinar, and Ümit V. Çatalyürek. Nucleus decompositions for identifying hierarchy of dense subgraphs. ACM Transactions on the Web (TWEB), 11(3):16:1– 16:??, July 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Shi:2020:TAW

[STZL20] Min Shi, Yufei Tang, Xingquan Zhu, and Jianxun Liu. Topicaware Web service representation learning. ACM Transactions on the Web (TWEB), 14 (2):9:1–9:23, April 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386041.

Singla:2011:CBC

[SW11] Adish Singla and Ingmar Weber. Camera brand congruence and camera model propagation in the Flickr social graph. ACM Transactions on the Web (TWEB), 5(4):20:1–20:??, October 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Su:2013:UQI

[SWL^{+}13] Weifeng Su, Hejun Wu, Yafei Li, Jing Zhao, Frederick H. Lochovsky, Hongmin Cai, and Tianqiang Huang. Understanding query interfaces by statistical parsing. ACM Transactions on the Web (TWEB), 7(2):8:1– 8:??, May 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Song:2016:IJV

[SXM^{+}16] Hengjie Song, Yonghui Xu, Huaqing Min, Qingyao Wu, Wei Wei, Jianshu Weng, Xiaogang Han, Qiang Yang, Jialiang Shi, Jiaqian Gu, Chunyan Miao, and Nishida Toyoaki. Individual judgments versus consensus: Estimating query-URL relevance. ACM Transactions on the Web (TWEB), 10(1):3:1– 3:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Shang:2023:CST

[SXY^{+}23] Zhihua Shang, Hongtao Xie, Lingyun Yu, Zhengjun Zha, and Yongdong Zhang. Constructing spatio-temporal graphs for face forgery detection. ACM Transactions on the Web (TWEB), 17(3):23:1–23:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580512.

Song:2011:ATR

[SZG11] Yang Song, Lu Zhang, and C. Lee Giles. Automatic tag recommendation algorithms for social recommender systems. ACM Transactions on the Web (TWEB), 5(1):4:1–4:??, February 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Sun:2015:ITB

[SZSA15] Chang-Ai Sun, Xin Zhang, Yan Shang, and Marco Aiello. Integrating transactions into BPEL service compositions: an aspectbased approach. ACM Transactions on the Web (TWEB), 9(2):9:1–9:??, May 2015. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Torre-Bastida:2018:RBT

[TBBI18] Ana I. Torre-Bastida, Jesús Bermúdez, and Arantza Illarramendi. A rule-based transducer for querying incompletely aligned datasets. ACM Transactions on the Web (TWEB), 12 (4):23:1–23:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Tonge:2020:IPP

[TC20] Ashwini Tonge and Cornelia Caragea. Image privacy prediction using deep neural networks. ACM Transactions on the Web (TWEB), 14(2): 7:1–7:32, April 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386082.

Tu:2018:ARP

[TCM^{+}18] Wenting Tu, David W. Cheung, Nikos Mamoulis, Min Yang, and Ziyu Lu. Activity recommendation with partners. ACM Transactions on the Web (TWEB), 12(1):4:1–4:??, February 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Tranquillini:2015:MEI

[TDKC15] Stefano Tranquillini, Florian Daniel, Pavel Kucherbaev, and Fabio Casati. Modeling, enacting, and integrating custom crowdsourcing processes. ACM Transactions on the Web (TWEB), 9(2):7:1–7:??, May 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Thomas:2014:UID

[Tho14] Paul Thomas. Using interaction data to explain difficulty navigating online. ACM Transactions on the Web (TWEB), 8(4):24:1–24:??, October 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Tuchinda:2011:BMD

[TKS11] Rattapoom Tuchinda, Craig A. Knoblock, and Pedro Szekely. Building mashups by demonstration. ACM Transactions on the Web (TWEB), 5(3):16:1– 16:??, July 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Tappenden:2009:CDS

[TM09] Andrew F. Tappenden and James Miller. Cookies: a deployment study and the testing implications. ACM Transactions on the Web (TWEB), 3(3):9:1–9:??, June 2009. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Thoma:2019:FEC

[TTHS19] Steffen Thoma, Andreas Thalhammer, Andreas Harth, and Rudi Studer. FusE: Entitycentric data fusion on linked data. ACM Transactions on the Web (TWEB), 13(2):8:1–8:??, April 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3306128.

Torres:2014:ASB

[TWH14] Sergio Duarte Torres, Ingmar Weber, and Djoerd Hiemstra. Analysis of search and browsing behavior of young users on the Web. ACM Transactions on the Web (TWEB), 8(2):7:1– 7:??, March 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Uribe:2017:UWP

[UÁM17] Silvia Uribe, Federico Álvarez, and José Manuel Menéndez. User’s Web page aesthetics opinion: a matter of lowlevel image descriptors based on MPEG-7. ACM Transactions on the Web (TWEB), 11(1):5:1– 5:??, March 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Umair:2024:EFA

[UBL^{+}24] Muhammad Umair, Athman Bouguettaya, Abdallah Lakhdari, Mourad Ouzzani, and Yuyun Liu. Exif2Vec: a framework to ascertain untrustworthy crowdsourced images using metadata. ACM Transactions on the Web (TWEB), 18(3):31:1– 31:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3645094.

Upadhyaya:2022:SFV

[UC22] Apoorva Upadhyaya and Joydeep Chandra. Spotting flares: The vital signs of the viral spread of tweets made during communal incidents. ACM Transactions on the Web (TWEB), 16(4):18:1– 18:??, November 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3550357.

Urvoy:2008:TWS

[UCFL08] Tanguy Urvoy, Emmanuel Chauveau, Pascal Filoche, and Thomas Lavergne. Tracking Web spam with HTML style similarities. ACM Transactions on the Web (TWEB), 2(1):3:1– 3:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Urgaonkar:2007:AMM

[UPS^{+}07] Bhuvan Urgaonkar, Giovanni Pacifici, Prashant Shenoy, Mike Spreitzer, and Asser Tantawi. Analytic modeling of multitier Internet applications. ACM Transactions on the Web (TWEB), 1(1):2:1–2:??, May 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Umyarov:2011:UEA

[UT11] Akhmed Umyarov and Alexander Tuzhilin. Using external aggregate ratings for improving individual recommendations. ACM Transactions on the Web (TWEB), 5(1):3:1– 3:??, February 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Uzun:2024:SRI

[Uzu24] Erdinç Uzun. Scraping relevant images from Web pages without download. ACM Transactions on the Web (TWEB), 18(1):1:1–1:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3616849.

Vahedian:2017:MRH

[VBM17] Fatemeh Vahedian, Robin Burke, and Bamshad Mobasher. Multirelational recommendation in heterogeneous networks. ACM Transactions on the Web (TWEB), 11(3):15:1– 15:??, July 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Vural:2014:SFW

[VCK14] A. Gural Vural, B. Barla Cambazoglu, and Pinar Karagoz. Sentiment-focused Web crawling. ACM Transactions on the Web (TWEB), 8(4):22:1–22:??, October 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Vassio:2018:YWY

[VDM^{+}18] Luca Vassio, Idilio Drago, Marco Mellia, Zied Ben Houidi, and Mohamed Lamine Lamali. You, the Web, and your device: Longitudinal characterization of browsing habits. ACM Transactions on the Web (TWEB), 12 (4):24:1–24:??, November 2018. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Vargiu:2013:ICA

[VGA13] Eloisa Vargiu, Alessandro Giuliani, and Giuliano Armano. Improving contextual advertising by adopting collaborative filtering. ACM Transactions on the Web (TWEB), 7(3):13:1–13:??, September 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Vlachos:2010:ODB

[VKY10] Michail Vlachos, Suleyman S. Kozat, and Philip S. Yu. Optimal distance bounds for fast search on compressed timeseries query logs. ACM Transactions on the Web (TWEB), 4 (2):6:1–6:??, April 2010. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Valderas:2011:SRS

[VP11] Pedro Valderas and Vicente Pelechano. A survey of requirements specification in modeldriven development of Web applications. ACM Transactions on the Web (TWEB), 5(2):10:1– 10:??, May 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Vidyapu:2021:IMW

[VVB21] Sandeep Vidyapu, Vijaya Saradhi Vedula, and Samit Bhattacharya. Investigating and modeling the Web elements’ visual feature influence on freeviewing attention. ACM Transactions on the Web (TWEB), 15(1):2:1–2:27, January 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3409474.

Victor:2013:ETB

[VVCD13] Patricia Victor, Nele Verbiest, Chris Cornelis, and Martine De Cock. Enhancing the trust-based recommendation process with explicit distrust. ACM Transactions on the Web (TWEB), 7(2):6:1–6:??, May 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Watanabe:2019:LCP

[WAP19] Willian Massami Watanabe, Giovana Lázaro Amêndola, and Fagner Christian Paes. Layout cross-platform and crossbrowser incompatibilities detection using classification of DOM elements. ACM Transactions on the Web (TWEB), 13(2):12:1–12:??, April 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3316808.

White:2008:LPD

[WBC08] Ryen W. White, Mikhail Bilenko, and Silviu Cucerzan. Leveraging popular destinations to enhance Web search interaction. ACM Transactions on the Web (TWEB), 2(3):16:1– 16:??, July 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Weerkamp:2012:EEC

[WBdR12] Wouter Weerkamp, Krisztian Balog, and Maarten de Rijke. Exploiting external collections for query expansion. ACM Transactions on the Web (TWEB), 6(4):18:1–18:??, November 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2023:ECR

[WCF^{+}23] Yingxu Wang, Xiaoru Chen, Jinyuan Fang, Zaiqiao Meng, and Shangsong Liang. Enhancing conversational recommendation systems with representation fusion. ACM Transactions on the Web (TWEB), 17(1):6:1–6:??, February 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3577034.

Wang:2015:SWU

[WCZ15] Yazhe Wang, Jamie Callan, and Baihua Zheng. Should we use the sample? Analyzing datasets sampled from Twitter’s stream API. ACM Transactions on the Web (TWEB), 9(3):13:1– 13:??, June 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Watanabe:2015:FQM

[WDD15] Willian Massami Watanabe, Ana Luiza Dias, and Renata Pontin De Mattos Fortes. Fona: Quantitative metric to measure focus navigation on rich Internet applications. ACM Transactions on the Web (TWEB), 9(4):20:1–20:??, October 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Watanabe:2023:LCB

[WdSdO23] Willian Massami Watanabe, Danilo Alves dos Santos, and Claiton de Oliveira. Layout cross-browser failure classification for mobile responsive design Web applications: Combining classification models using feature selection. ACM Transactions on the Web (TWEB), 17 (4):32:1–32:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580518.

Wu:2024:CCG

[WGF24] Huizi Wu, Cong Geng, and Hui Fang. Causality and correlation graph modeling for effective and explainable session-based recommendation. ACM Transactions on the Web (TWEB), 18(1):2:1–2:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593313.

White:2013:CBD

[WH13] Ryen W. White and Eric Horvitz. Captions and biases in diagnostic search. ACM Transactions on the Web (TWEB), 7(4):23:1–23:??, October 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

White:2014:CBO

[WH14] Ryen W. White and Ahmed Hassan. Content bias in online health search. ACM Transactions on the Web (TWEB), 8(4):25:1–25:??, October 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Walk:2016:ADC

[WHGS16] Simon Walk, Denis Helic, Florian Geigl, and Markus Strohmaier. Activity dynamics in collaboration networks. ACM Transactions on the Web (TWEB), 10(2):11:1– 11:??, May 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

White:2021:WMN

[Whi21] Ryen W. White. Welcome message from the new Editorin-Chief. ACM Transactions on the Web (TWEB), 15(3): 11e:1, July 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3456294.

Wu:2013:MVC

[WHS13] Ou Wu, Weiming Hu, and Lei Shi. Measuring the visual complexities of Web pages. ACM Transactions on the Web (TWEB), 7(1):1:1–1:??, March 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Weninger:2013:PPF

[WJH13] Tim Weninger, Thomas J.

Johnston, and Jiawei Han. The parallel path framework for entity discovery on the web. ACM Transactions on the Web (TWEB), 7(3):16:1–16:??, September 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wu:2020:SAR

[WLCG20] Zhiang Wu, Changsheng Li, Jie Cao, and Yong Ge. On scalability of association-rule-based recommendation: a unified distributed-computing framework. ACM Transactions on the Web (TWEB), 14(3): 13:1–13:21, July 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3398202.

Wang:2023:HGT

[WLP^{+}23] Shuhai Wang, Xin Liu, Xiao Pan, Hanjie Xu, and Mingrui Liu. Heterogeneous graph transformer for meta-structure learning with application in text classification. ACM Transactions on the Web (TWEB), 17(3):21:1–21:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3580508.

Wang:2021:ACN

[WLT^{+}21] Wei Wang, Jiaying Liu, Tao Tang, Suppawong Tuarob, Feng Xia, Zhiguo Gong, and Irwin King. Attributed collaboration network embedding for academic relationship mining. ACM Transactions on the Web (TWEB), 15(1):4:1– 4:20, January 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3409736.

Wu:2023:DBC

[WML23] Bin Wu, Zaiqiao Meng, and Shangsong Liang. Dynamic Bayesian contrastive predictive coding model for personalized product search. ACM Transactions on the Web (TWEB), 17 (4):33:1–33:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3609225.

Wu:2016:QDQ

[WMS^{+}16] Wensheng Wu, Weiyi Meng, Weifeng Su, Guangyou Zhou, and Yao-Yi Chiang. Q2P: Discovering query templates via autocompletion. ACM Transactions on the Web (TWEB), 10 (2):10:1–10:??, May 2016. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2023:CPS

[WMTO23] Xiao Wang, Craig MacDonald, Nicola Tonellotto, and Iadh Ounis. ColBERT-PRF: Semantic pseudo-relevance feedback for dense passage and document retrieval. ACM Transactions on the Web (TWEB), 17(1):3:1–3:??, February 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3572405.

Weinreich:2008:QAE

[WOHM08] Harald Weinreich, Hartmut Obendorf, Eelco Herder, and Matthias Mayer. Not quite the average: an empirical study of Web use. ACM Transactions on the Web (TWEB), 2(1):5:1– 5:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2022:BBS

[WOM22] Xi Wang, Iadh Ounis, and Craig Macdonald. BanditProp: Bandit selection of review properties for effective recommendation. ACM Transactions on the Web (TWEB), 16(4):20:1– 20:??, November 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532859.

Weber:2013:FBW

[WPB13] Ingo Weber, Hye-Young Paik, and Boualem Benatallah. Formbased Web service composition for domain experts.

ACM Transactions on the Web (TWEB), 8(1):2:1–2:??, December 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2022:LSE

[WPC^{+}22] Kai Wang, Jun Pang, Dingjie Chen, Yu Zhao, Dapeng Huang, Chen Chen, and Weili Han. A large-scale empirical analysis of ransomware activities in bitcoin. ACM Transactions on the Web (TWEB), 16 (2):7:1–7:29, May 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3494557.

Wang:2021:IEA

[WQG^{+}21] Huan Wang, Chunming Qiao, Xuan Guo, Lei Fang, Ying Sha, and Zhiguo Gong. Identifying and evaluating anomalous structural change-based nodes in generalized dynamic social networks. ACM Transactions on the Web (TWEB), 15(4): 19:1–19:22, July 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3457906.

Wilson:2019:APP

[WSL^{+}19] Shomir Wilson, Florian Schaub, Frederick Liu, Kanthashree Mysore Sathyendra, Daniel Smullen, Sebastian Zimmeck, Rohan Ramanath, Peter Story, Fei Liu, Norman Sadeh, and Noah A. Smith. Analyzing privacy policies at scale: From crowdsourcing to automated annotations. ACM Transactions on the Web (TWEB), 13(1):1:1–1:??, February 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wilson:2012:BSG

[WSPZ12] Christo Wilson, Alessandra Sala, Krishna P. N. Puttaswamy, and Ben Y. Zhao. Beyond social graphs: User interactions in online social networks and their implications. ACM Transactions on the Web (TWEB), 6(4):17:1–17:??, November 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wu:2011:TDQ

[WST11] Mingfang Wu, Falk Scholer, and Andrew Turpin. Topic distillation with query-dependent link connections and page characteristics. ACM Transactions on the Web (TWEB), 5(2):6:1– 6:??, May 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2024:DAG

[WWC^{+}24] Pengfei Wang, Daqing Wu, Chong Chen, Kunpeng Liu, Yanjie Fu, Jianqiang Huang, Yuanchun Zhou, Jianfeng Zhan, and Xiansheng Hua. Deep adaptive graph clustering via von Mises–Fisher distributions. ACM Transactions on the Web (TWEB), 18(2):22:1– 22:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3580521.

Wu:2020:STI

[WWN^{+}20] Huijun Wu, Chen Wang, Richard Nock, Wei Wang, Jie Yin, Kai Lu, and Liming Zhu. SMINT: Toward interpretable and robust model sharing for deep neural networks. ACM Transactions on the Web (TWEB), 14(3): 11:1–11:28, July 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381833.

Wang:2017:VMC

[WWW^{+}17] Tianyi Wang, Gang Wang, Bolun Wang, Divya Sambasivan, Zengbin Zhang, Xing Li, Haitao Zheng, and Ben Y. Zhao. Value and misinformation in collaborative investing platforms. ACM Transactions on the Web (TWEB), 11(2):8:1– 8:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2018:OWP

[WYJ^{+}18] Yue Wang, Dawei Yin, Luo Jie, Pengyuan Wang, Makoto Yamada, Yi Chang, and Qiaozhu Mei. Optimizing wholepage presentation for Web search. ACM Transactions on the Web (TWEB), 12(3):19:1– 19:??, July 2018. CODEN ????

ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2015:DCU

[WYY^{+}15] Jing Wang, Clement T. Yu, Philip S. Yu, Bing Liu, and Weiyi Meng. Diversionary comments under blog posts. ACM Transactions on the Web (TWEB), 9(4):18:1–18:??, October 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2024:CGS

[WZJ^{+}24] Luzhi Wang, Yizhen Zheng, Di Jin, Fuyi Li, Yongliang Qiao, and Shirui Pan. Contrastive graph similarity networks. ACM Transactions on the Web (TWEB), 18(2): 17:1–17:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580511.

Wang:2017:CUB

[WZT^{+}17] Gang Wang, Xinyi Zhang, Shiliang Tang, Christo Wilson, Haitao Zheng, and Ben Y. Zhao. Clickstream user behavior models. ACM Transactions on the Web (TWEB), 11(4):21:1– 21:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Wang:2016:STQ

[WZZ^{+}16] Xinyu Wang, Jianke Zhu, Zibin Zheng, Wenjie Song, Yuanhong Shen, and Michael R. Lyu.

A spatial-temporal QoS prediction approach for time-aware Web service recommendation. ACM Transactions on the Web (TWEB), 10(1):7:1–7:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Xiao:2024:PPI

[XJZL24] Chunjing Xiao, Wanlin Ji, Yuxiang Zhang, and Shenkai Lv. PIDKG: Propagating interaction influence on the dynamic knowledge graph for recommendation. ACM Transactions on the Web (TWEB), 18(2): 26:1–26:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593314.

Xiao:2020:PRF

[XLC20] Zhijun Xiao, Cuiping Li, and Hong Chen. PatternRank+NN: a ranking framework bringing user behaviors into entity set expansion from Web search queries. ACM Transactions on the Web (TWEB), 14(3): 10:1–10:15, July 2020. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386042.

Xiao:2009:BSD

[XLH^{+}09] Xiangye Xiao, Qiong Luo, Dan Hong, Hongbo Fu, Xing Xie, and Wei-Ying Ma. Browsing on small displays by transforming Web pages into hierarchically structured subpages. ACM Transactions on the Web (TWEB), 3(1):4:1–4:??, January 2009. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Xiao:2010:LSS

[XLL^{+}10] Xiangye Xiao, Qiong Luo, Zhisheng Li, Xing Xie, and Wei-Ying Ma. A large-scale study on map search logs. ACM Transactions on the Web (TWEB), 4(3):8:1–8:??, July 2010. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Xu:2017:EIE

[XLWS17] Haitao Xu, Daiping Liu, Haining Wang, and Angelos Stavrou. An empirical investigation of ecommerce-reputation-escalation-as-a-service. ACM Transactions on the Web (TWEB), 11(2):13:1–13:??, May 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Xiao:2023:DKE

[XVWK23] Yunming Xiao, Matteo Varvello, Marc Warrior, and Aleksandar Kuzmanovic. Decoding the Kodi ecosystem. ACM Transactions on the Web (TWEB), 17(1):2:1–2:??, February 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563700.

Xu:2017:COF

[XZ17] Chang Xu and Jie Zhang. Collusive opinion fraud detection in online reviews: a probabilistic modeling approach. ACM Transactions on the Web (TWEB), 11(4):25:1– 25:??, September 2017. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yu:2008:FWS

[YB08] Qi Yu and Athman Bouguettaya. Framework for Web service query algebra and optimization. ACM Transactions on the Web (TWEB), 2(1):6:1– 6:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yang:2008:DGN

[YL08] Bo Yang and Jiming Liu. Discovering global network communities based on local centralities. ACM Transactions on the Web (TWEB), 2(1):9:1– 9:??, February 2008. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yu:2019:EPP

[YMZ19] Weiren Yu, Julie McCann, and Chengyuan Zhang. Efficient pairwise penetrating-rank similarity retrieval. ACM Transactions on the Web (TWEB), 13 (4):21:1–21:??, December 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yue:2013:MSI

[YW13] Chuan Yue and Haining Wang.

A measurement study of insecure JavaScript practices on the Web. ACM Transactions on the Web (TWEB), 7(2):7:1– 7:??, May 2013. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yang:2023:RAS

[YYL^{+}23] Yingguang Yang, Renyu Yang, Yangyang Li, Kai Cui, Zhiqin Yang, Yue Wang, Jie Xu, and Haiyong Xie. RoS-GAS: Adaptive social bot detection with reinforced selfsupervised GNN architecture search. ACM Transactions on the Web (TWEB), 17(3):15:1– 15:??, August 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3572403.

Yu:2024:BNB

[YZD^{+}24] Fangyu Yu, Peng Zhang, Xianghua Ding, Tun Lu, and Ning Gu. BNoteHelper: a note-based outline generation tool for structured learning on video-sharing platforms. ACM Transactions on the Web (TWEB), 18(2): 28:1–28:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3638775.

Yu:2007:EAW

[YZL07] Tao Yu, Yue Zhang, and Kwei-Jay Lin. Efficient algorithms for Web services selection with end-to-end QoS constraints. ACM Transactions on the Web (TWEB), 1(1):6:1–6:??, May 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Yu:2024:LNU

[YZZ^{+}24] Mei Yu, Kun Zhu, Mankun Zhao, Jian Yu, Tianyi Xu, Di Jin, Xuewei Li, and Ruiguo Yu. Learning neighbor user intention on user-item interaction graphs for better sequential recommendation. ACM Transactions on the Web (TWEB), 18 (2):21:1–21:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580520.

Zhang:2021:STR

[ZB21] Shuo Zhang and Krisztian Balog. Semantic table retrieval using keyword and table queries. ACM Transactions on the Web (TWEB), 15(3): 11:1–11:33, July 2021. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3441690.

Zafar:2015:SCO

[ZBG^{+}15] Muhammad Bilal Zafar, Parantapa Bhattacharya, Niloy Ganguly, Krishna P. Gummadi, and Saptarshi Ghosh. Sampling content from online social networks: Comparing random vs. expert sampling of the Twitter stream. ACM Transactions on the Web (TWEB), 9(3):12:1– 12:??, June 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zhao:2023:DGM

[ZC23] Chenye Zhao and Cornelia Caragea. Deep gated multimodal fusion for image privacy prediction. ACM Transactions on the Web (TWEB), 17 (4):34:1–34:??, November 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3608446.

Zhu:2023:MTG

[ZCC^{+}23] Guixiang Zhu, Jie Cao, Lei Chen, Youquan Wang, Zhan Bu, Shuxin Yang, Jianqing Wu, and Zhiping Wang. A multi-task graph neural network with variational graph auto-encoders for session-based travel packages recommendation. ACM Transactions on the Web (TWEB), 17(3):18:1– 18:??, August 2023. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3577032.

Zheng:2010:UTM

[ZCL^{+}10] Yu Zheng, Yukun Chen, Quannan Li, Xing Xie, and Wei-Ying Ma. Understanding transportation modes based on GPS data for Web applications. ACM Transactions on the Web (TWEB), 4(1):1:1–1:??, January 2010. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zdun:2007:MPD

[ZHD07] Uwe Zdun, Carsten Hentrich, and Schahram Dustdar. Modeling process-driven and serviceoriented architectures using patterns and pattern primitives. ACM Transactions on the Web (TWEB), 1(3):14:1–14:??, September 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zubiaga:2020:EDS

[ZJ20] Arkaitz Zubiaga and Aiqi Jiang. Early detection of social media hoaxes at scale. ACM Transactions on the Web (TWEB), 14 (4):18:1–18:23, September 2020. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3407194.

Zhang:2021:EWD

[ZJZ^{+}21] Jifeng Zhang, Wenjun Jiang, Jinrui Zhang, Jie Wu, and Guojun Wang. Exploring weather data to predict activity attendance in event-based social network: From the organizer’s view. ACM Transactions on the Web (TWEB), 15 (2):10:1–10:25, May 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3440134.

Zhang:2021:SUC

[ZLD^{+}21] Peng Zhang, Baoxi Liu, Xianghua Ding, Tun Lu, Hansu Gu, and Ning Gu. Studying and understanding characteristics of post-syncing practice and goal in social network sites. ACM Transactions on the Web (TWEB), 15(4): 16:1–16:26, July 2021. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3457986.

Zhang:2024:HGN

[ZLG^{+}24] Guangping Zhang, Dongsheng Li, Hansu Gu, Tun Lu, and Ning Gu. Heterogeneous graph neural network with personalized and adaptive diversity for news recommendation. ACM Transactions on the Web (TWEB), 18(3):34:1– 34:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3649886.

Zhao:2019:USE

[ZLKL19] Liping Zhao, Pericles Loucopoulos, Evangelia Kavakli, and Keletso J. Letsholo. User studies on end-user service composition: a literature review and a design framework. ACM Transactions on the Web (TWEB), 13(3):15:1–15:??, October 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3340294.

Zou:2023:PTL

[ZLL^{+}23] Lixin Zou, Weixue Lu, Yiding Liu, Hengyi Cai, Xiaokai Chu, Dehong Ma, Daiting Shi, Yu Sun, Zhicong Cheng, Simiu Gu, Shuaiqiang Wang, and Dawei Yin. Pre-trained language model-based retrieval and ranking for Web search. ACM Transactions on the Web (TWEB), 17(1):4:1–4:??, February 2023. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3568681.

Zhao:2022:CAD

[ZLW^{+}22] Yiji Zhao, Youfang Lin, Zhihao Wu, Yang Wang, and Haomin Wen. Context-aware distance measures for dynamic networks. ACM Transactions on the Web (TWEB), 16(1): 2:1–2:34, February 2022. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3476228.

Zhang:2016:DSP

[ZLZL16] Xianchao Zhang, Zhaoxing Li, Shaoping Zhu, and Wenxin Liang. Detecting spam and promoting campaigns in Twitter. ACM Transactions on the Web (TWEB), 10(1):4:1–4:??, February 2016. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zhu:2012:CLS

[ZM12] Guangyu Zhu and Gilad Mishne.

ClickRank: Learning sessioncontext models to enrich Web search ranking. ACM Transactions on the Web (TWEB), 6 (1):1:1–1:??, March 2012. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zareie:2024:FIM

[ZS24] Ahmad Zareie and Rizos Sakellariou. Fuzzy influence maximization in social networks. ACM Transactions on the Web (TWEB), 18(3):32:1–32:??, August 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3650179.

Zeginis:2011:CDR

[ZTC11] Dimitris Zeginis, Yannis Tzitzikas, and Vassilis Christophides. On computing deltas of RDF/S knowledge bases. ACM Transactions on the Web (TWEB), 5 (3):14:1–14:??, July 2011. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zheng:2024:HIC

[ZWH^{+}24] Xiaolin Zheng, Rui Wu, Zhongxuan Han, Chaochao Chen, Linxun Chen, and Bing Han. Heterogeneous information crossing on graphs for sessionbased recommender systems. ACM Transactions on the Web (TWEB), 18(2):19:1– 19:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3572407.

Zhang:2014:PBT

[ZWML14] Xianchao Zhang, You Wang, Nan Mou, and Wenxin Liang. Propagating both trust and distrust with target differentiation for combating link-based Web spam. ACM Transactions on the Web (TWEB), 8(3):15:1– 15:??, June 2014. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zhang:2015:REA

[ZWZL15] Haibin Zhang, Yan Wang, Xiuzhen Zhang, and Ee-Peng Lim. ReputationPro: The efficient approaches to contextual transaction trust computation in E-commerce environments. ACM Transactions on the Web (TWEB), 9(1):2:1– 2:??, January 2015. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zhu:2024:ITG

[ZXC^{+}24] Haorui Zhu, Fei Xiong, Hongshu Chen, Xi Xiong, and Liang Wang. Incorporating a triple graph neural network with multiple implicit feedback for social recommendation. ACM Transactions on the Web (TWEB), 18 (2):23:1–23:??, May 2024. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580517.

Zheng:2016:PQA

[ZYZ16] Huiyuan Zheng, Jian Yang, and Weiliang Zhao. Probabilistic QoS aggregations for service composition. ACM Transactions on the Web (TWEB), 10 (2):12:1–12:??, May 2016. CO-DEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zhang:2024:DAF

[ZZL^{+}24] Peng Zhang, Qi Zhou, Tun Lu, Hansu Gu, and Ning Gu. DeLink: an adversarial framework for defending against cross-site user identity linkage. ACM Transactions on the Web (TWEB), 18(2):29:1– 29:??, May 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

https://dl.acm.org/doi/10.
1145/3643828.

Zheng:2011:RFL

[ZZM^{+}11] Yu Zheng, Lizhu Zhang, Zhengxin Ma, Xing Xie, and Wei-Ying Ma. Recommending friends and locations based on individual location history. ACM Transactions on the Web (TWEB), 5(1):5:1–5:??, February 2011. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Zhang:2024:SSE

[ZZW^{+}24] Lei Zhang, Wuji Zhang, Likang Wu, Ming He, and Hongke Zhao. SHGCN: Socially enhanced heterogeneous graph convolutional network for multibehavior prediction. ACM Transactions on the Web (TWEB), 18(1):4:1–4:??, February 2024. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617510.